A Bibliography of The Mathematica Journal

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 0.10

Title word cross-reference

1/π [Bag13]. 2 [Tro99c]. 3 [Ful96b, SB94a, SB94b]. n [Fre93a]. p [De 99]. π [Abb05b].

-adic [De 99]. -Queens [Fre93a].

1 [RM13, TiS96].

2.0 [Lev92c]. 2.1 [Wei94].

3.0 [Ano97c, Lic96, Pet96, Sof96, Str96, Zel97]. 3000 [Phi97b].

4 [Ger97, GG97, Kna97]. 4.2 [Hoh99].

5 [The99]. 5.0 [Ner99].

Abstract [Mae92b]. access [Ano06]. Accessing [Gay94a, KG90]. Accuracy [Ner99]. Achieving [MG10]. Acoustic [CGW11, RMWP05]. Ad [Cha94]. adic [De 99]. Aerodynamics [Fea05]. Affine [McC06]. Affymetrix [All13]. against [Vrb11a]. Agent [dO10]. Aided [Mil90, Mur96, SYS^{+}90]. Airfoil [Fea05]. Algebra [AMP97, Mat98, Nag05, Vrb12b]. Algebraic [Koz99, MD99, SYS^{+}90, Str97, Ziz92b, Ziz92a]. Algebriac [Str96]. Algorithm [Car94, Fre93b, Jon06, Str94]. Algorithmic [Bar06]. Algorithms [Mae94a, Mae96c]. Amplitude [Sko94]. Amusement [Rud95b]. Analysis [Abb99c, Car95, CP90, DEH92, DHE92, FPP95, HLF96, KP99, Kor93c, Mal14, Mil90, Rid90, Wag96b, Yel10]. Analytica [CZ93]. Analytical [HL94]. Anamorphograms [Kuc99]. Andrei [Far92]. Angles [Sar99].

Animated [Mae94a, SMH98]. Animating [ST90]. Animation [Mur92]. Animations [Mur94]. Announcements [Ano94e, Ano94f, Lev92d]. ANOVA [CC11]. Answer [Abb97e, Abb97f, Abb98c, Abb98d, Abb98e, Abb99d, Abb99e, Abb99f, Abb05c]. Application [DEH92, DHE92, EP06, KMRC90]. Applications [AUV06, EMM90, Fre94, JG98, Mae94c, Wei99]. Approach [Bar06, BM99, Sil14, SYS^{+}90]. Approaching [Aiz97]. Approximants [Pro99]. Approximate [CSA94]. Approximating [Nev99]. Approximation [Abb99a, CD92b]. Approximations [CD92a]. Arbelos [RM14]. Area [Ano97a]. Arithmetic [De 99, Kei95, Mae96b]. Arrays [Kna97, Tro06]. Arrows [Ful96a]. Art [Phi97a, Sko92]. Artificial [Fre93a, Fre93b, Fre94]. Artlandia [Phi97a]. Asked [Chr93a, Zel97]. Asymptotic [Abb05b, Coh95c, Win12]. August [Abb05a, Ker98]. Australia [Abb05a]. Austria [Ker98]. AuthorTools [San99a]. Automata [dOV14]. Automatic [LO90]. available [Ano06]. Axis [RM14].

Backer [Fre94]. Backer-Upper [Fre94]. Bag [Chr92b, Chr92a, Chr93a, Chr93b, wDJK92]. Banach [Wag93]. Barcodes [Abb99c]. Based [Buc98, DSM05, KJ06, NS95, Pro99, Sel94]. Bases [Hel95, Lic96, NG94]. Basic [Bel90a, Bel90b, Fre93b, Mae96b, Mar11]. Basis [Kel06]. Bayesian [Cas14, Kor93a]. Beauty [Ent99]. Behavior [CJ98, McP92b, McP92a]. Bessel [Aul12]. Betting [Vrb11a]. Beyond [Peg06b, Peg05a, Peg05b, Peg05c, Peg06a]. Bifurcation [Mur94]. Binary [Gay94a]. Binomial [Zwi13]. BioEqCalc [AG98]. Black [Mil90]. Blocks [Jac92d]. Board [Pap05]. Bodies [Mur92]. Book [Cha94, Mae95a, Mat95, Wei94]. Books [Smi93]. Boolean [Cow13]. Bootstrap [Var99]. Boundaries [McC06]. Boundary [PFS06]. Bravais [MM93]. Brick [Cou13]. Brownian [Mae96a]. Browser [San99b]. Browsers [Har94b]. Buckets [RM06]. Buckminsterfullerenes [MM93]. Buffon [Joh06, Sin06]. BUGS [Mae95a]. Building [Gay99]. Built [Cou13, Sof93]. Built-in [Sof93]. Buttons [Mil97, Nov96a].

C [FES05]. Calculating [Coh95a]. Calculation [San96]. Calculator [Gre98, Bri99]. Calculus [Bal94, BPU90, Ede92a, HR92]. Calendar [Ano94a, Ano94b, Ano95a]. Canonical [Mal14]. Capital [Kor94a]. Case [ST90]. Cauchy [Vrb11b]. Cellular [dOV14]. CFLP [MIS98]. challenges [And90]. Change [dO10]. Chaos [Mae95c, Sha95]. Chapter [Wel13]. Characters [Sko94]. Charged [Vrb12b]. Chemistry [Ram97]. Chicago [Bel92]. Chocolate [MNOH13]. Circle [RM14]. Circuit [Rid90]. Classes [Mae98]. Classic [Peg05a]. Classroom [Ano97c, Coh95a, Coh95b, Coh96, Par14, Ram97]. Classrooom [Coh95c]. CleanSlate [Gay93a]. Clearing [Gay93a]. Clipping [Hel98]. Closure [MG10]. Clusters [KBS12]. Code [FES05, JG98, Mae93c, Var96]. Coefficients [Nev99]. Coin [HN96]. Color [Ful94a, Mat98]. Combinatorics [Ker06]. Combinators [Mae95b]. Compass [Hel92]. Competition [Ano92, Lev92a]. Compiler [Liu96]. Complete [Ede92c, Ede92b]. Complex [Abb97b, Ano97d, EP06, Mar98, Sha95, Sil14]. ComplexExpand [Gay94c]. Complexity [Sut06]. Compositions [Tro98a, Tro98b]. Compression [Wat94]. Computation [AS95, Cow13, Mae98, Phi98a, Wol05]. Computational [Lya99, RM13, RS99]. Computations [Kei92, Mae90]. Compute [Par14]. Computer [BM94, CP90, Mat98, Mil90, Mur96, Phi97a, SYS^{+}90]. Computer-Aided [Mil90, Mur96, SYS^{+}90]. Computers [Gun90]. Computing [Abb97b, Abb99c, Agu90, BCDH94, CC11, Chr92a, Her06, NG94]. Concerning [Mar11]. Conchyallometry [Sto95]. Concrete [Lya05]. Conditioned [KMRC90]. Cone [Edw97]. Configuration [KP99]. Confluent [AS95]. Conjecture [Ede92c, Ede92b]. Conjugates [Abb97b]. Constants [Tro05a]. Constraint [MIS98]. Constructing [Bag13, Tro06]. Construction [MD99]. Constructions [Hel92]. Contexts [Wag96a]. Continued [Ada92b, Ada92a, Tro99b]. Continuous [Aul12, Hra12, Win12]. Contour [WJ93]. Control [Fre94, JG98, Wit96]. Controlling [Cou13]. Conundrums [Ano98]. Conversion [Wit92]. Convolution [Knu92]. Cool [Phi97a]. Coordinates [SB94b]. Corner [Tro97a, Tro97b, Tro98a, Tro98b, Tro98c, Tro99b, Tro99a, Tro99d, Tro99c, Tro05a, Tro05b, Tro06]. Correlation [Mal14]. Correspondence [Yel10]. Cosine [Wat94]. Courseware [BPU90]. Court [Cha05]. Cover [Kab98]. Creating [Dil92a]. Critical [Jon06, Mit96]. Crossing [Peg06a]. Crossword [Tro06]. Crystal [End97]. Crystals [MJ93]. Cumulant [MG10]. Curriculum [FW92]. Curve [Ful94a, Rud06]. Curves [ABBS10, Cas14, Mae92d, RR14]. Customizing [Gay94b, Mil97, San99b]. Cylinder [Coh96]. Cylindrical [Aul12, Hra12, Win12].

D [Ful96b, SB94a, SB94b, Tro99c]. Data [Bla92a, Bla93a, Ful96b, Mae92b, Mae96b, Nov99, Pal14, Smi90a, Wit90]. Database [vGU99]. Databases [Mae93b]. Debugging [Mae95a]. Decisions [Kor93c, Kor93d, Kor93a, Kor93b, Kor94a, Kor94b]. Definite [Ada96]. Degree [Jon06]. Demographics [Wei92]. Demonstrations [Peg05a]. Demystifying [Bla98]. Densities [Vin99]. Density [Pro99, Rei98]. Dependency [Wag96b]. Design [Ano97b, CC11, KP99, PFS06]. designed [Héb97]. Designer [Kuz94, Kuz95a, Kuz95b]. Detecting [All13]. Determinants [Chr92a]. Development [BM99, RLK05]. Diagrams [Mar98]. Dialogs [Mil97]. Differences [SF11]. Different [Pal14]. Differential [All13, Ano98, Boc93, Her99, Ker12, Kna96, Lya99, Nev99, Pet92, Ziz92b, Ziz92a]. Differentiation [Abb98b]. Diffraction [Sen94]. Diffusion [Hra12]. Dimensional [Bla93b, SMH98, Sul92, WJ93]. Directed [Hu99]. Discovered [MNOH13]. Discrete [AUV06, CD92b, CD92a, Wat94]. Discretely [RN05]. Discrimination [MHT11]. Disease [DSM05]. Disordered [Coh95a]. Displaying [End97]. Distributed [MIS98]. Distribution [Ent99, Rei98, RS96, Thi95, Vrb11b]. Distribution-Free [Thi95]. Distributions [Abb97c, RN99]. Dither [tHR05]. Divisibility [Sut06]. DNA [Abb99c]. Document [San99a]. Documents [Dil92b]. Does [Sko92]. Doing [Ner99]. Domain [Aul12]. Domains [MD99]. Dominoes [Rud95c]. Dot [BN92, Tro99c]. Dots [MW05]. Drag [Gre98]. Drag-and-Drop [Gre98]. Drawing [Hu99]. Drawings [Héb90]. DRIMA [dO10]. Drop [Gre98]. Drum [Tro96, Tro99d]. Dynamic [Lya05, Wag95]. Dynamical [HL94, Jon06]. Dynamics [Aul12, Hra12, Ker12, Win12, dO10].

E-mail

[Chr92b, Chr92a, Chr93a, Chr93b, wDJK92]. Each [Vrb11a]. Earth [SMH98]. Econometrics [Bel90a, Bel90b]. Econometrics.m [Bel90b, Bel90a]. Economics [Vin99]. Economy [Nog92b].

Edge [Hel92]. Editing [Kuz95a]. Editor [Phi97a, Phi98a]. Education [Bal94, HR92, PUS92, SMH98]. Educational [ARAM92]. Effect [Peg05c]. Efficient [Hay95, Hu99, Mae96c]. Eigenfunctions [Bon12]. Eight [Cha94]. Einstein [Agu90]. Electron [MW05]. Electronic [Chr92b, MW05]. Elegant [Var92]. Element [Dai98, DT98]. Elementary [Tro98a, Tro98b, Wat90]. Elements [BM99]. Elimination [Ede92c, Ede92b]. Ellipse [Abb06, Tro96]. Ellipse-Shaped [Tro96]. Elliptic [LT99]. Elsewhere [Kuz95b]. Embedded [Tro99c]. Employing [Mar98]. End [Gay93c, Gay94b, Gra92, Gra96, NS95]. end.m [Chr93b]. Engineer [Jac92d, Jac92c, Jac92a, Jac92b]. Engineering [HLF96, RLK05]. Enhanced [Gay93b]. Enhancement [HR92]. Ensemble [Vin99]. Entanglement [Nor95]. Environment [Fri99, RLK05]. Epidemic [GVV92]. Equation [CGI95, CGW11, ST90]. Equations [Abb99a, Abb99b, Ano98, Boc93, Her99, HL94, Kam05, Ker12, Kna96, Lya99, Nev99, Pet92, Ric99, Tro99b]. Equilibria [DK92a, DK92b, Par14]. Equilibrium [Nog92a, Nog92b]. Equivalence [Mae98]. Error [Gay93b, Pal14, Zel96]. Escher [LO90]. Estimating [Rud06]. Estimators [Vrb11b]. Euclidean [RM13]. Euler [PUS92]. European [RN05]. Evaluation [Kel06, PFS06]. Events [Ano94a, Ano94b, Ano95a]. everyone [Ano06]. Evolutionary [KJ06]. Exact [Nev99]. Example [Thi95, Vrb11b, Kor93a]. Examples [Lya99, Sel94]. Excel [Ful94a]. Exception [Mae97c]. Excerpts [PUS92]. Exercises [AMP97]. Exotic [RN05]. Expanding [CGW11]. Expansion [Abb05b]. Expansions [Coh95c, Nak93]. Expectations [Ken99]. Experiments [Hay95, Jak96]. Experts [Abb97e, Abb97f, Abb98c, Abb98d, Abb98e, Abb99d, Abb99e, Abb99f, Abb05c]. Explicit [Abb99b]. Exploratory [KJ06]. Exploring [Cas14, Hra12, Pap05, RR14, Win12]. Exponents [San96]. Exporting [Bla92a]. Exports [Ful94b]. Expression [All13]. Expressions [Har94a, Mez92]. Extended [Rud06]. Extending [Sof93]. Extensions [Mae94d]. Exterior [Ede92a]. External [KG90]. Extracting [Nov99].

Faces [Gra90]. Facial [DT98]. Facing [Smi90a]. Factorizations [KM99]. False [Ede92c, Ede92b]. Families [RM99, dOV14]. Fast [Mae92c, Mae98, Rei92, AUV06]. Faster [Tro06]. Features [Mae96d]. Fibonacci [JR98, Mae92c, RR14, Rei92]. Files [Gay94a, Abb98b]. Filtering [Rud06]. Finance [Lya99]. Financial [Kel06, Mil90, Vin99]. Finding [DK92a, DK92b, Jon06, Tro05a]. Finite [Dai98, DT98, SF11, Sut92b, Sut92a]. First [FW92, Vrb12a]. First-Year [FW92]. Fisher [MHT11]. Fits [Ful94a]. Fitting [Bla92c, Pal14]. Five [Bla93d]. Flexible [NBT99]. Flipping [Kor93b]. Floating [Jac92a]. Flow [Coh96]. Fluid [Coh96]. Fonts [Ano97b]. Force [Hu99]. Force-Directed [Hu99]. Formal [Koe94, Nak93]. Format [Sof93, Ano06]. Formatted [Abb97c]. Forms [Ede92a, Ziz92b, Ziz92a]. Formulas [Bag13, MNOH13]. Fortran [FES05]. Forum [Bal94, HR92, PUS92]. Foundations [Wol98]. Four [McP92b, McP92a, Sul92]. Four-Dimensional [Sul92]. Four-Layer [McP92b, McP92a]. Fourier [AUV06, Abb98b, Aul12, Hra12, Win12]. Fourier-Space [Aul12, Hra12, Win12]. Fractal [Cas14, Far92, GIRR97, Mae92d, RR14]. Fractals [Kor93c]. Fraction [Tro99b]. Fractional [Mae96a]. Fractions [Ada92b, Ada92a]. Frames [DT98]. Framework [RN05]. Free [Abb05b, Ano06, Thi95]. Freeing [Gay93a]. Frequently [Chr93a, Zel97]. Front [Gay93c, Gay94b, Gra92, Gra96, NS95]. Function [AS95, Abb99b, Har94b, Her06, Mae95c, MF12]. Functional [CGI95, Wag97]. Functionals [CD92a]. Functions [Ano97d, Aul12, CD92b, EP06, Kel06, LT99, Lya05, Mae95d, Rei98, Tro98a, Tro98b, Tro99a, Wei99, Wit96]. Future [Wol05, Gun90]. Fuzzy [Fre94].

G [TiS96]. Gained [Kor94a]. Gambler [Vrb12a]. Game [Peg06b, Pap05, Par14]. Games [Cou93, MNOH13]. Garnet [MJ93]. Gauss [Abb99a, Edi90]. Gaussian [Ede92c, Ede92b, Wag96d]. Genaille [Thi93]. Gene [All13]. General [Bag13, Nog92a, Nog92b]. Generalizations [RR14, RM14]. Generalized [Abb99a]. Generate [Mik13]. Generating [Abb99c, Gay94b, GIRR97, McC06, Sul92, Wei99]. Generation [FES05, HL94, JG98]. Generator [Fie99]. Genetic [Fre93b, Nac95]. Geode [Ebn90]. Geographic [vGU99]. Geometric [Cou95]. Geometry [AD99, Aul12, Bec93, Hra12, RM13, Win12]. Ghost [Cha94, Lev96]. Glacial [SYS^{+}90]. Global [Mae97a, Mae97b, Wei92]. Golf [Mur93]. Google [Abb05b]. Graph [Hu99, Mat98, Sil14]. Graphic [Kuz95a]. Graphical [Nog92b]. Graphics [Ano92, BM94, Bla93b, Ful94b, Kuz94, Kuz95a, Kuz95b, Lev92b, Mae94d, Phi90, Smi92a, SB94a, SB94b]. Graphing [Ben05]. Graphs [Mar11, Peg06a]. Gröbner [Hel95, Lic96, NG94]. Group [PC99].

Hadamard [Abb05b]. Hagenberg [Ker98]. Handled [Dic92]. Handling [Gay94b, Mae97c]. Hare [Wei94]. Hashing [Jac92b]. Heat [ST90]. Helicatenoid [Oga92a]. Help [Har94b, San99b]. Here [Mat95]. High [BTC96, Hu99, Jon06, KMRC90, MNOH13]. High-Degree [Jon06]. High-Precision [KMRC90]. High-Quality [BTC96, Hu99]. Higher [Mae95d]. Higher-Order [Mae95d]. History [Chr93b]. Holaphics [Jak94]. Holes [Hor05]. Hyperbolic [Wag93]. Hypergeometric [AS95]. Hypothesis [Aiz97].

Icosian [Peg06b]. If [Kor94b]. II [Mae94c]. Ill [KMRC90]. Ill-Conditioned [KMRC90]. Illustrating [Smi92b]. Image [GG97, Jan98, Kab98, Mae95e, Wat94]. Images [GIRR97]. Imaging [KMRC90]. Implementation [Jon06, Mil90, NBT99, RMWP05]. Implementing [BSS93, Cas14, Gre98]. Implicit [SNG92]. Importing [Bla92a, Bla93a]. Imports [Ful94b]. Improving [Jac92c]. Incircle [RM14]. Including [Mez92]. Inclusion [MJ93]. Independence [AD99]. Index [Ano95b]. Individual [DSM05]. Individual-Based [DSM05]. Industrial [BM99]. Inequalities [Abb99b, Str94, Str97]. Inference [Cas14, Kor93a, Vin99]. Infinite [Aul12, Mae90]. Infinitum [Cha94]. Infinity [Cha94]. Inflight [Ano98]. Information [AD99]. Insects [DSM05]. Installation [Gay93c]. Integer [BC96, Kor93d]. Integers [KM99, Mae96b, Mae96c, NG94]. Integral [Ano98, Aul12, Lya99, Ric99]. Integrals [Bon12, Ken99]. Integrated [FW92, RLK05]. Integrating [Not96]. Integration [Ada96, Coh95c, Ful96b, Lya05, Saa96]. Intelligence [Fre93a, Fre93b, Fre94]. Interactive [AMP97, Dil92b, Phi90, RK05]. InterCall [Wei94]. Interface [CW92b, CW92a, KG90, Phi90]. Interfaces [Gay99]. Interior [SMH98]. International [Abb05a, Ker98]. Interpolants [MD99]. Interpolating [Lya05]. Interpolation [Abb98b, Abb99a]. Interpretation [Wag93]. Interpreter [Mae94b]. Interval [Kei95, Mae97a, Mae97b]. Introduction [Abb05a, Ker98, Mat95, Wel13, Yel10]. Inversion [CSA94, KMRC90, Kuc99]. Inversive [RM13]. Involving [Bon12]. issues [Ano06]. Iterates [Nev99]. Iteration [Mae95c, Wag97].

J [Gay99]. J/Link [Gay99]. Journal [Ano06]. Jumping [RM06]. Jurisprudence [Cha05]. Just [Bla93d].

Kalman [Rud06]. Kepler [Vrb11c]. Kernels [Gay94c, MHT11]. Kerr [SR90]. Kim [Jon92a]. Kind [Hra12, Win12]. Kinetic [Bel92, Mar98]. Kit [Dub90a, Dub90b]. Knot [Gof95]. Knowledge [Bla92b]. Kuhn [Abb99a, Kam05].

Labels [Ful94b, Lev92b]. Laboratories [HR92]. Ladders [Edw94]. Lagrange [Edw94]. Lamp [Rud95a]. Landing [Mur12]. Laplace [CSA94]. Large [Gay94b]. Lattices [Coh95a, MM93]. Lava [Rud95a]. Law [CJ98]. Layer [McP92b, McP92a]. Lazy [Cha94]. Learning [Bla92a, Bla92b, Bla93a, Bla93b, Bla93c, CJ98, Her99, RK05]. Legendre [Abb99a]. Legends [Ful95a]. LEGO(R) [Cou13]. Leibniz [Gre98]. Length [Coh95b]. Lévy [RN05]. Libraries [KG90]. Lighting [SB94a]. Limits [Ada92b, Ada92a, Phi98b]. Linde [Far92]. Linear [Car94, Car95, CD92b, CD92a, Jan98, Kor93d, Ner99, Nev99]. Liners [Abb90a, Abb90b, AL92a, AL92b, Jou92a]. Link [Peg05b, Gay99]. Lists [Gay94b]. Local [Ful96a]. Logic [Coh92b, Coh92a, Mae94b, Mae94c, SN06, Wat90]. Logistic [Mik13]. Long [Mae96b, Mae96c]. Looping [Bla92c]. Luge [Rud98]. Lyapunov [San96].

M [TiS96]. M/G/1 [TiS96]. Machine [Her98, Cha94]. Machines [Mae93c, NBT99, Sut92b, Sut92a]. Macintosh [Gra92]. Magic [Sar99]. mail [Chr92b, Chr92a, Chr93a, Chr93b, wDJK92]. Making [Hor05, Kuz94, Wei05]. Manifolds [Bar06]. Manipulating [Hut95, SA90]. Manipulation [GG97]. Manipulators [HL94]. Many [Dic92]. Many-Handled [Dic92]. Map [Mik13]. Mapping [Wes93]. Marks [Ful95b]. Matching [Fal05, Har97a]. MathCode [FES05]. Mathematica [Abb05a, Ano06, Ker98, Abb90a, Abb90b, Abb05b, ARAM92, Aiz97, AH92, And90, Ano92, Ano94c, Ano94d, Ano95c, Ano97b, Bal94, Bla92a, Bla92b, Bla93a, Bla93b, Bla93c, BPU90, Buc98, BC96, CW92b, CW92a, Car94, Car95, CZ93, Cou93, Foa97, FES05, Gay93a, Gay93b, Gay94a, Gay95, Ger97, GG97, Har97a, Héb97, Hoh99, Jac92a, Jac92b, Jon06, Jon92a, Jon92b, KP99, Kei95, Kel06, Kna97, Koz99, Kuz95a, Lev92c, Lev92b, Lic96, Mae90, Mae92c, Mae92d, Mae92e, Mae92g, Mae92b, Mae92f, Mae93a, Mae93b, Mae93c, Mae93d, Mae94a, Mae94b, Mae94c, Mae94d, Mae95b, Mae95c, Mae95d, Mae95e, Mae96a, Mae96b, Mae96c, Mae97c, Mae97b, Mat95, MNOH13, Mur12, Ner99, Not96, Pet96, Phi90, Phi97b, Ram97, RN99, Rud95b, San99c, SB93, Sin06, Sof96, Str96, The99, Tro97a]. Mathematica [Wel13, Wol98, Jou92a, Mat95]. Mathematica-designed [Héb97]. Mathematical [Fri99, Mur93, Soi96, Tro99a]. Mathematics [Wol98, Gun90]. MathLink [BTC96, Bor92, Ful94a, Ful95a, Gay94a, NS95, Wag96c]. MathLink-Based [NS95]. MathModelica [Fri99]. MathPDE [SF11]. MathTensor [CP90]. MathWorld [Wei05]. Matrices [Abb99a, Coh95b, KMRC90]. Matrix [Ano98, AMP97]. Maxillo [DT98]. Maxillo-Facial [DT98]. Mechanical [DEH92, DHE92, HL94]. Mechanism [Erd06]. Media [Aul12, Hra12, SR90, Win12]. Medical [KMRC90]. Memoization [Cas14]. Memory [Gay93a, Gay94c]. Message [Mae92f]. Messages [Gay93b, Zel96]. Metamorphic [MJ93]. Method [Bag13, Dai98, Edi90, Fal05]. Methods [Coh95a, Fea05, Lya99, Win12]. Microarrays [All13]. MINDSTORMS(R) [Cou13]. Minimal [BCDH94, Dic90, Mae92e, dO10]. Mir [Foa97]. Mixed [CC11, Kor93d]. Mixed-Design [CC11]. Mixed-Integer [Kor93d]. Mixing [SR90]. ML [Vrb11b]. Moats [Edw94, Wag96d]. Möbius [Gof95, Rud95c]. Modal [McP92b, McP92a]. Mode [Fal05, Wag96c]. Mode-Matching [Fal05]. Model [AH92, Mil90, Nog92a, Nog92b]. Modeling [Fri99, Hra12, Kor94a, SNG92]. Models [BSS93, DSM05, Pal14]. Modular [PC99, Tro99b]. Modulo [Ent99]. Molecular [DEH92, DHE92]. Moment [MG10, Pro99]. Moment-Based [Pro99]. Moments [TiS96]. Monitored [RN05]. Moon [Mur12]. Morphological [Jan98]. Morphology [End97]. Motion [Bec93, Mae96a, Sar99, Sar11, Vrb12b]. Mu [Rud06]. Multi [MW05, dO10]. Multi-Agent [dO10]. Multi-Electron [MW05]. Multichannel [Fal05]. Multiple [Hra12, Hut95]. Multipliers [Edw94]. Multivariate [Nak93, RS96]. Musical [JR98]. Must [Kor94b].

Nash [DK92a, DK92b, Par14]. Needle [Joh06, Sin06]. Negative [Zwi13]. Neglect [MG10]. Neighbourhoods [AD99]. Nested [Ada92b, Ada92a]. net [And90]. Nets [MM93]. Network [BCDH94, Cha05, Fre93a]. Neural [And90, Fre93a]. News [Ano94c, Ano94d, Ano95c]. Newton [Edi90, Edw97]. Nine [Bla93d]. Nineties [BPU90]. Nodal [Rid90]. nohup [Chr93b]. Noise [Abb99a]. Non [RM13]. Non-Euclidean [RM13]. Nondecreasing [Abb99b]. Nonlinear [Aul12, Bla92c, Boc93, Edi90, Hra12, Jon06, Ner99, Sar11, Win12]. Nonparametric [Cas14, Wei99]. Normal [RS96]. Nose [Edw97]. Nose-Cone [Edw97]. Notebook [BTC96, Dub90a, Dub90b, Mez92]. Notebooks [Nov96b, Smi90b, Smi93]. Notes [Coh95a, Coh95b, Coh95c, Coh96]. Nothing [Kor94a]. Number [Peg05b, Peg06a]. Numbers [BTC96, Rei92, Str96]. Numeric [Nag05]. Numerical [Abb98b, BSS93, Ful95b, Ful96b, Kei92, KG90, Ner99, Nog92b, Saa96, San96]. Numerics [Sof96]. NXT [Cou13].

Object [BM99, Fri99, Mae93a].

Object-Oriented [BM99, Fri99, Mae93a]. ODEs [Win12]. Offshore [PFS06]. One [Abb90a, Abb90b, AL92a, AL92b, Ent99, Jou92a]. One-Liners [Abb90a, Abb90b, AL92a, AL92b, Jou92a]. online [Ano06]. online-only [Ano06]. only [Ano06]. Opening [Smi90b]. Operator [Ano98]. Optimal [Lya05]. Optimization [KP99, KJ06, Mae93c, Mae97a, Mae97b, Saa96, ST90]. Option [BSS93]. Options [Kel06, RN05]. Order [Mae95d, RS99]. Ordered [KM99]. Ordinary [Nev99]. Oriented [BM99, Fri99, Mae93a]. Orloj [Geh97]. Orthogonal [Rei98]. Other [Vrb11a, Abb97b]. Overview [Lev92c, Pet96].

Package [Bel90a, Bel90b, End97, Ful95c, Gay93a, Koe94, NS95, San99a, SF11, Wat90]. Packages [Bla93c, Coh92b, Coh92a, Smi90b, Dil92a]. Packed [Kna97]. Padé [Abb99a]. Palette [Abb98b]. Palettes [Nov96a]. Panel [Fea05]. Paperless [Ano06]. Parabolic [Lya99]. Paradigms [Hra12]. Paradox [Cou93, Wag93]. ParentForm [Abb99c]. Park [Rud95b]. Part [Mae97b, RM13]. Partial [Ker12, Kna96, Lya99]. Particle [Vrb12b]. Pascal [Abb99a]. Passage [Vrb12a]. Passing [Mae92f]. Past [Coh96]. Path [Lya99]. Pattern [Har97a]. Patterns [Har94a, Vrb11a]. Pauli [Vrb12b]. PDEs [SF11]. Pedagogical [Joh06]. Mathematica [Dil92a, Dil92b]. Pendulum [Sar11]. Perfect [Kuz94]. Performance [Hay92b, Hay92a, Jac92c]. Performing [CP90]. Perimeter [Abb06]. Periodic [Nev99]. Perspective [Joh06]. Perspectives [SB94b]. Perth [Abb05a]. Perturbation [DEH92, DHE92]. Pets [RM06]. Phase [Mur94, Sko94]. Pi [Coh95a]. Picard [Nev99]. Pick [Phi97a, Phi98a]. Picture [Kuz94]. Pictures [Cou95, Kuz95b]. Pipes [Jac92d]. Pivoting [Ede92c, Ede92b]. Placing [KBS12]. Planar [McP92b, McP92a, Mur94]. Plane [Sha95]. Playing [Cha94]. Plot [CC11]. Plots [Gay94b, Mur94, TC94, WJ93]. Plotting [Mae97a, Mae97b, Wes93]. Point [Jac92a]. Polar [Tow96, TC94]. Polygonal [MD99]. Polygons [Hel98, SA90]. Polyhedra [KBS12, Mae93d]. Polymorphism [Mae92f]. Polynomial [Str94]. Polynomials [Fie99, Hut95, Jon06, Knu92, Nag05]. Polyominoes [RM99]. Polytopes [Sul92]. Population [Ker12]. possible [Gun90]. Post [SYS^{+}90]. Post-Glacial [SYS^{+}90]. PostScript [Chr92b]. Potent [Cou93]. Power [Cou93, Koe94, Nak93, Tan97, Wag95, Wag96a, Wag96b, Wag96c, Wag97]. Powers [Bag13, Nak93]. Practice [ARAM92]. Precision [Ful95b, KMRC90]. Prediction [FPP95]. Presentations [EP06]. Pricing [BSS93, RN05]. Printed [Sko94]. Probabilistic [Cas14]. Probability [Rei98, SB93]. Probing [dO10]. Problem [Edw97, FW92, Fre93a, SYS^{+}90, Vrb11c, Joh06]. Problems [Cow13, Fal05, Lya05, PFS06]. Procedural [Wit96]. Proceedings [Ker98]. Process [RN05]. Processes [AD99]. Processing [EMM90, Jan98, San99a, Wit90]. Produced [KBS12]. Product [Aul12]. Profiler [Gay95]. Profiles [Jon92a]. Program [Ano94e, Ano94f, BTC96, DK92a, DK92b, Gay94a, Lev92d, Phi90, Rid90, Wit92]. Programmer [Mae90, Mae92c, Mae92d, Mae92e, Mae92g, Mae92b, Mae92f, Mae93a, Mae93b, Mae93c, Mae93d, Mae94a, Mae94b, Mae94c, Mae94d, Mae95b, Mae95c, Mae95d, Mae95e, Mae96a, Mae96b, Mae96c, Mae97c, Mae97b]. Programming [BC96, Car94, Car95, Cas14, Ful95c, Hay95, Hra12, Kor93d, Koz99, Lev92a, Mae92g, Mae93a, Mae94b, Mae94c, Mae96d, Mat95, Nac95, Nov96b, Sel94, Wag95, Wag96a, Wag96b, Wag96c, Wag97, Wel13, Ziz92b, Ziz92a, Abb97b]. Programs [Hay92b, Hay92a, Jac92c, Var92]. Projectile [Sar99]. Projection [SB94b]. Prolog [Not96]. Proofs [Abb99b, RM14]. Propagator [RMWP05]. Properties [RR14, RM13]. Prose [Mae95a]. Protein [FPP95]. Proteins [JR98]. Prototyping [Erd06]. Prover [CZ93]. Proving [Buc98]. Pseudo [Var96]. Putting [Kuz95b]. Puzzles [Peg05a, Tro97a].

Quadrature [Abb99a]. Quality

[BTC96, Hu99]. Quantum [DEH92, DHE92, Her98, MW05, Phi98a, Tro99c]. Quasicrystals [Sen94]. Quasirandom [Car11]. Quaternions [PT98]. Queens [Fre93a]. Questions [Abb97e, Abb97f, Abb98c, Abb98d, Abb98e, Abb99d, Abb99e, Abb99f, Abb05c, Chr93a, Ful95a, Zel97]. Queueing [AH92]. Queues [TiS96].

Radial [Kel06]. Radical [RM14]. Radicals [Ada92b, Ada92a]. Rado [Her06]. Ramanujan [Bag13, Tro99b]. Ramanujan-Type [Bag13]. Ramsey [Cow13]. Random [Ano97a, AD99, BM94, BN92, BTC96, Coh95b, Mar11, Sil13]. Random-Dot [BN92]. Randomness [Phi98b]. Rank [Mit96]. Rational [Coh95b, LT99]. Rational-Length [Coh95b]. Ray [Mae94d]. Reactive [dO10]. ReadList [Ful94b]. Real [PT98]. Rearranging [Har94a]. Reason [Sko92]. Reasons [Sko92]. Rebound [SYS^{+}90]. Reconstructing [MJ93]. Recursion [Bla92c]. Reduce [Abb99a, Kam05, Par14]. Reflective [Kuc99]. Regal [Lev96]. Region [RMWP05]. Regression [Edi90, Ner99, Zwi13]. Regular [AS95, Bla93d]. Regularization [Abb05b]. Related [RM99]. Relational [Mae93b]. Relativistic [Vrb12b]. Reliability [Sil14]. Remote [Gay94c]. Removal [tHR05]. Rendering [GIRR97, Jak96, Sul92]. Replacements [Erd06]. Representation [Aul12, Hra12, Rei98, Win12]. Representations [JR98, Nog92b]. Representing [GVV92, dOV14]. Resampling [SB93]. Research [SMH98]. Results [Mar11]. Return [Tro05b]. Review [Cha94, Har94b, Hoh99, Mae95a, Mat95, Wei94]. Revisited [Peg06b, Hay92b, Hay92a, Joh06, Oga92b]. Rhombic [KBS12]. Ricci [Agu90]. Ride [Rud98]. Riemann [Aiz97, Ben05, Kiv06, Tro97b, Tro98c, Tro05b]. Rigid [Mur92]. Ripple [Peg05c]. Robotica [NS95]. Robots [Cou13]. Rods [Thi93]. Rogers [Tro99b]. Rook [Fie99]. Roots [Abb99b]. Rotating [Mur92]. Rotational [DEH92, DHE92]. Rotations [PT98]. Ruin [Vrb12a]. Rule [Sel94]. Rule-Based [Sel94]. Rules [Bla98, Sof93, dOV14].

S [CW92b, CW92a]. Sachs [CGW11]. Sampler [Var92]. Sampling [Vrb11b]. Sand [Héb97]. Scattering [Fal05, Tro99c]. Scholes [Mil90]. School [MNOH13]. Scoping [Mae92a]. Scott [Jon92a]. Scratching [Mik13]. Search [Abb05b].

Searching [Tro99a]. Second [Hra12, Sar99, Win12]. Sector [Nog92a, Nog92b]. Seismic [Nov99]. Selected [RM13]. Self [McC06]. Self-Affine [McC06]. Semantic [Har97a]. Semantica [Har97a]. Semigroups [Sut92b, Sut92a]. Sensitivity [Car95]. Sequences [Abb99c]. Series [BM94, Cou95, Koe94, Kor93c, Nak93, Rei98]. Set [Abb05b, Coh92b, Coh92a, Sar99]. Setting [dO10]. Seven [Gra90]. Seventhg [Abb05a]. Shading [SB94a]. Shadowing [Wag96a]. Shaped [Tro96]. Shapes [DT98]. Sigma [Her06]. Signal [EMM90]. Signed [Mit96]. SimCity [Phi97b]. Simple [Sar11]. Simplex [Car94]. Simulate [Kor94b]. Simulating [Fre93b]. Simulation [Car11, Fri99, JG98]. Simulations [Ram97]. Simulator [Her98, Bel92]. Sin [Gra90]. Single [Mae95e]. Single-Image [Mae95e]. Singularities [Hra12, Win12]. Site [Tro99a]. Skew [Vin99]. Slicing [Moo95]. Slinky [Rud94]. Slip [Rud06]. Smooth [Abb98b, MD99]. snapshots [Gun90]. Soft [Mur12]. Software [BM99, Har94b]. Solid [SNG92]. Solids [Bla93d, Moo95]. Solution [Fre93a]. Solutions [Nev99]. Solve [Abb99a, Cow13, Kam05, SF11]. Solvers [Boc93]. Solving [AH92, CGI95, FW92, Str97, Tro97a, Vrb11c, Wes93]. Some [AUV06, Bon12, Cow13, Lya05, Sel94]. Sound [Coo92, Tro99d]. Sounds [Mik13]. Space [Aul12, Hra12, Win12]. Spatial [FW92, Kuc99]. Sphere [Ben05]. Spider [Jon06]. Spinors [Nor95]. Split [CC11, RMWP05]. Split-Plot [CC11]. Spread [DSM05, GVV92]. Stable [RN99]. Stacks [Har94b]. Standard [Bla93c]. Start [Mat95]. State [Sut92b, Sut92a, Sut06]. States [Tro99c]. Statistic [Mit96]. Statistical [Abb97c]. Statistics [RS99, SB93]. Stats [SB93]. Stereograms [BN92, Mae95e]. Stochastic [Cas14, Ken99]. Stopping [Lya05]. Straight [Hel92].

Strategies [Pap05]. String [JR98]. Strings [Wel13]. Strong [Str94]. Structure [Cha05, FPP95, MW05, SMH98]. Structures [Mae90]. Structuring [ARAM92]. Students [MNOH13]. Study [Sar11, ST90]. Subclass [Dub90a, Dub90b]. Subexpressions [Cop92]. Substitutions [Erd06]. Sudoku [Peg06b, Peg05a, Peg05b, Peg05c, Peg06a]. Sum [Abb05b]. Sum-Free [Abb05b]. Super [Sar11]. Super-Nonlinear [Sar11]. Support [Ful94a, Ful94b, Ful95a, Ful95b, Ful95c, Ful96a, Ful96b, Gay93c, Gay94b, Gay94c, Mil97, NBT99, Zel96, Zel97]. Supreme [Cha05]. Surface [Gay94b, MM93, SN06, Tro05b]. Surfaces [ABBS10, BCDH94, Dic90, Dic92, HLF96, Hor05, Kiv06, Mae92e, MJ93, Tro97b, Tro98c]. Survey [Wit90]. Swarm [KJ06]. Swarm-Based [KJ06]. Swing [Mur93]. Symbolic [Abb99c, Boc93, Coh95c, EMM90, Nag05, PFS06]. Symbolic-Numeric [Nag05]. Symbols [Gay93a, Mae92a]. Symmetric [Bon12, Sha95]. Symmetric-Top [Bon12]. Symmetries [Her99]. Symmetry [TC94]. Symposium [Abb05a, Ker98]. Syntactic [Sut92b, Sut92a]. System [Buc98, CP90, FES05, Mae92g, MIS98, Sil14, dO10]. Systems [Fre94, Jon06, Str94, Tan97].

Tables [Abb97c]. Taking [Bal94]. Tarski [Wag93]. Tax [Abb99c]. Teaching [Jon92b]. Tech [Ful94a, Ful94b, Ful95a, Ful95b, Ful95c, Ful96a, Ful96b, Gay93c, Gay94b, Gay94c, Zel96, Zel97]. Technical [Hoh99, Mil97]. techniques [Abb97b]. Technology [PFS06]. Tensor [CP90]. Tensors [Agu90]. Test [Bla92b]. Testing [Thi95]. Tests [Wei99]. TEX/Mathematica [Dil92a, Dil92b]. Their [Her99, Ken99, McC06]. Themes [RM13]. Theorem [CZ93, CGW11]. Theorema [Buc98]. Theory [ARAM92, Coh92b, Coh92a, Cow13, DEH92, DHE92, Jon06, Par14, Sil14]. There [Bla93d]. Third [Ker98]. Three [Aul12, PUS92, SMH98, WJ93]. Three-Dimensional [SMH98, WJ93]. Throwing [Sin06]. Thumbtacks [Kor93b]. Tick [Ful95b]. Tiles [McC06]. Tilings [Sen94]. Time [Kor93c, Vrb12a]. Tomography [Mur96]. Tool [Dub90a, Dub90b, GVV92]. Toolbox [Car11, Gay93a, Gay93b, Gay94a, Gay95, Jac92d, Jac92c, Jac92a, Jac92b]. Toolkit [KJ06]. Tools [KP99, Par14, Phi97a]. Top [Bon12]. Topics [Chr92b]. Topographies [HN96]. Tortoise [Wei94]. Towers [Cou93]. Tracing [Mae94d]. Tracings [Héb97]. Track [Mae92c]. Trade [AL92c, Abb92a, Abb92b, Abb92c, Abb93a, Abb93b, Abb93c, Abb93d, Abb94a, Abb94b, Abb94c, Abb94d, Abb95a, Abb95b, Abb95c, Abb95d, Abb96a, Abb96b, Abb96c, Abb96d, Abb97a, Abb97b, Abb97c, Abb98a, Abb98b, Abb99a, Abb99b, Abb99c, Abb05b, Ano98, LA92, Jou92b]. Transcendental [Abb99b, Tro98a, Tro98b]. Transform [CSA94, Wat94, AUV06]. Transforms [EMM90]. Transient [SR90]. Transputer [BCDH94]. Trefoil [Gof95]. Triacontahedra [KBS12]. Triangle [Ano97a]. Tricks [AL92c, Abb92a, Abb92b, Abb92c, Abb93a, Abb93b, Abb93c, Abb93d, Abb94a, Abb94b, Abb94c, Abb94d, Abb95a, Abb95b, Abb95c, Abb95d, Abb96a, Abb96b, Abb96c, Abb96d, Abb97a, Abb97b, Abb97c, Abb98a, Abb98b, Abb99a, Abb99b, Abb99c, Abb05b, Ano98, Kam05, LA92, Jou92b]. Trinoid [Oga92b]. Trott [Tro97a, Tro97b, Tro98a, Tro98b, Tro98c, Tro99b, Tro99a, Tro99d, Tro99c, Tro05a, Tro05b, Tro06]. Truck [Fre94]. Tubes [Edw93]. Tucker [Abb99a, Kam05]. Turing [Cha94, Her98, Mae93c]. Tutorial [SNG92, Var99]. Tutorials [ARAM92]. Twins [RM14]. Twisted [Edw93]. Two [Bla93b, Mae97b, Mar11, Nog92a, Nog92b, SR90, Vrb11a]. Two-Dimensional [Bla93b]. Two-Sector [Nog92a, Nog92b]. Two-Wave [SR90]. Type [Bag13]. Types [Mae92b, Mae96b]. Typesetting [Soi96]. Tzitzeica [ABBS10].

Uncertainty [Kor93c, Kor93d, Kor93a, Kor93b, Kor94a, Kor94b, Phi98b]. Uncomputable [Her06]. Understand [Sko92]. Uniform [Ent99, Mae93d]. Universe [CGW11, Far92]. UNIX [Gay93c]. unlimited [Ano06]. Unordered [KM99]. Unusual [Tro99d]. Updating [vGU99]. Upper [Fre94]. Use [Gay94c]. User [Gay99]. Using [Abb98b, Abb99a, ARAM92, All13, Bla93c, Cow13, Fea05, Gay99, KP99, Kam05, Kel06, Mik13, MNOH13, Par14, Phi90, Wat94].

Values [Mae92a, Mit96]. Variables [Ful96a, Hut95]. Vector [NBT99]. Vectors [Coh95b, Ede92a]. Venture [Kor94a]. Ventured [Kor94a]. Version [Ger97, Ano97c, Zel97]. Vertices [KBS12]. Very [KMRC90]. via [AD99, Jon06, Rud06, Smi93]. Vibrating [Tro96]. Vibrational [DEH92, DHE92]. Vibrational-Rotational [DEH92, DHE92]. Virtual [Rud94]. Visualization [Ano97d, Ful96b, GVV92, Kiv06, Tro97b, Tro98c]. Visualizing [EP06]. Volume [Ano95b]. Voting [Tan97].

Walker [MF12]. Walks [Mar11, Sil13]. Warning [Zel96]. Waterworks [Rud97]. Wave [CGW11, RMWP05, SR90]. Waveguide [McP92b, McP92a, Tro99c]. Wavelets [Nov99, RR95]. WaveX [Nov99]. Web [Sil13]. Weeds [DSM05]. Weighted [Tan97]. Western [Abb05a]. Wide [Sil13]. Wilcoxon [Mit96]. Windows [Hor05, Gay94c]. Within [MJ93]. without [RM14]. Wolfe [CGW11]. Wolfram [Tro99a, Peg05a]. Wooden [Gof95]. Word [RR14]. Words [Ker06, RM14]. work [And90]. Working [Cop92]. World [Sil13]. WorldPlot [Nov93]. Writing [Mae95a].

X [Gay93c, Gay94b]. Xmath [Bri99]. XML [San99c].

Year [FW92].

Zero [Cha94]. Zeros [Jon06]. Zonohedra [Tow96]. Zonohedrification [Har97b].

References

Abbott:1990:MOLa

[Abb90a] Paul C. Abbott. Mathematica one-liners. Mathematica Journal, 1(1):??, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                                  //www.mathematica-journal.
                                                  com/issue/v1i1/tutorials/oneliners/
                                                  index.html.

Abbott:1990:MOLb

[Abb90b] Paul C. Abbott. Mathematica one-liners. Mathematica Journal, 1(2):38–39, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                                  //www.mathematica-journal.
                                                  com/issue/v1i2/tutorials/oneliners/
                                                  index.html.

Abbott:1990:IOa

[Abb90c] Paul C. Abbott. In[] and Out[]. Mathematica Journal, 1(1):??, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                                  //www.mathematica-journal.

com/issue/v1i1/tutorials/inout/
index.html.

Abbott:1990:IOb

[Abb90d] Paul C. Abbott. In[] and Out[]. Mathematica Journal, 1 (2):??, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/tutorials/inout/
index.html.

Abbott:1991:TTb

[Abb92a] Paul Abbott. Tricks of the trade. Mathematica Journal, 2 (2):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i2/tutorials/tricks/
index.html.

Abbott:1992:TTa

[Abb92b] Paul Abbott. Tricks of the trade. Mathematica Journal, 2(3): ??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i3/tutorials/tricks/
index.html.

Abbott:1992:TTb

[Abb92c] Paul Abbott. Tricks of the trade. Mathematica Journal, 2 (4):??, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i4/tutorials/tricks/
index.html.

Abbott:1991:IOc

[Abb92d] Paul Abbott. In[] and Out[]. Mathematica Journal, 2(2):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i2/tutorials/inout/
                                        index.html.

Abbott:1992:IOa

[Abb92e] Paul Abbott. In[] and Out[]. Mathematica Journal, 2(3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i3/tutorials/inout/
                                        index.html.

Abbott:1992:IOb

[Abb92f] Paul Abbott. In[] and Out[]. Mathematica Journal, 2(4):??, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i4/tutorials/inout/
                                        index.html.

Abbott:1991:IOb

[Abb92g] Paul C. Abbott. In[] and Out[]. Mathematica Journal, 2 (1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/tutorials/inout/
                                        index.html.

Abbott:1993:TTa

[Abb93a] Paul Abbott. Tricks of the trade. Mathematica Journal, 3(1):

??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i1/columns/tricks/
index.html.

Abbott:1993:TTb

[Abb93b] Paul Abbott. Tricks of the trade. Mathematica Journal, 3 (2):??, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i2/tutorials/tricks/
index.html.

Abbott:1993:TTc

[Abb93c] Paul Abbott. Tricks of the trade. Mathematica Journal, 3(3): ??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/tutorials/tricks/
index.html.

Abbott:1993:TTd

[Abb93d] Paul Abbott. Tricks of the trade. Mathematica Journal, 3 (4):??, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i4/tutorials/tricks/
index.html.

Abbott:1993:IOa

[Abb93e] Paul Abbott. In[] and Out[].

Mathematica Journal, 3(1):??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.

                                        com/issue/v3i1/columns/inout/
                                        index.html.

Abbott:1993:IOb

[Abb93f] Paul Abbott. In[] and Out[]. Mathematica Journal, 3(2):??, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v3i2/tutorials/inout/
                                        index.html.

Abbott:1993:IOc

[Abb93g] Paul Abbott. In[] and Out[]. Mathematica Journal, 3(3):??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v3i3/tutorials/inout/
                                        index.html.

Abbott:1993:IOd

[Abb93h] Paul Abbott. In[] and Out[]. Mathematica Journal, 3(4):??, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v3i4/tutorials/inout/
                                        index.html.

Abbott:1994:TTa

[Abb94a] Paul Abbott. Tricks of the trade. Mathematica Journal, 4(1):38– 42, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i1/tutorials/tricks/
                                        38-42_tricks41.mj.pdf;   http:
                                        //www.mathematica-journal.

com/issue/v4i1/tutorials/tricks/
index.html.

Abbott:1994:TTb

[Abb94b] Paul Abbott. Tricks of the trade. Mathematica Journal, 4 (2):??, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i2/tutorials/tricks/
index.html.

Abbott:1994:TTc

[Abb94c] Paul Abbott. Tricks of the trade. Mathematica Journal, 4(3): ??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i3/tutorials/tricks/
index.html.

Abbott:1994:TTd

[Abb94d] Paul Abbott. Tricks of the trade. Mathematica Journal, 4 (4):??, Fall 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i4/tutorials/tricks/
index.html.

Abbott:1994:IOa

[Abb94e] Paul Abbott. In[] and Out[].

Mathematica Journal, 4(1):28– 37, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i1/tutorials/inout/
28-37_InOut41.mj.pdf;   http:
//www.mathematica-journal.

                                        com/issue/v4i1/tutorials/inout/
                                        index.html.

Abbott:1994:IOb

[Abb94f] Paul Abbott. In[] and Out[]. Mathematica Journal, 4(2):??, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i2/tutorials/inout/
                                        index.html.

Abbott:1994:IOc

[Abb94g] Paul Abbott. In[] and Out[]. Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i3/tutorials/inout/
                                        index.html.

Abbott:1994:IOd

[Abb94h] Paul Abbott. In[] and Out[]. Mathematica Journal, 4(4):??, Fall 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i4/tutorials/inout/
                                        index.html.

Abbott:1995:TTa

[Abb95a] Paul Abbott. Tricks of the trade. Mathematica Journal, 5(1): ??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i1/tutorials/tricks/
                                        index.html.

Abbott:1995:TTb

[Abb95b] Paul Abbott. Tricks of the trade. Mathematica Journal, 5 (2):??, Spring 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i2/tutorials/tricks/
index.html.

Abbott:1995:TTc

[Abb95c] Paul Abbott. Tricks of the trade. Mathematica Journal, 5(3): ??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i3/tutorials/tricks/
index.html.

Abbott:1995:TTd

[Abb95d] Paul Abbott. Tricks of the trade.

Mathematica Journal, 5(4):27– 34, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/tutorials/tricks/
27-34trix.54.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v5i4/tutorials/tricks/
index.html.

Abbott:1995:IOa

[Abb95e] Paul Abbott. In[] and Out[].

Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i1/tutorials/inout/
index.html.

Abbott:1995:IOb

[Abb95f] Paul Abbott. In[] and Out[]. Mathematica Journal, 5(2):??, Spring 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i2/tutorials/inout/
                                        index.html.

Abbott:1995:IOc

[Abb95g] Paul Abbott. In[] and Out[]. Mathematica Journal, 5(3):??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i3/tutorials/inout/
                                        index.html.

Abbott:1995:IOd

[Abb95h] Paul Abbott. In[] and Out[]. Mathematica Journal, 5(4):22– 26, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i4/tutorials/inout/
                                        22-26inout54.mj.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v5i4/tutorials/inout/
                                        index.html.

Abbott:1996:TTa

[Abb96a] Paul Abbott. Tricks of the trade. Mathematica Journal, 6(1): ??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i1/tutorials/tricks/
                                        index.html;   http://www.mathematica-

journal.com/issue/v6i1/tutorials/
tricks/tricks61.nb.

Abbott:1996:TTb

[Abb96b] Paul Abbott. Tricks of the trade. Mathematica Journal, 6 (2):??, Spring 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i2/tutorials/tricks/
index.html;   http://www.mathematica-
journal.com/issue/v6i2/tutorials/
tricks/tricks62.nb.

Abbott:1996:TTc

[Abb96c] Paul Abbott. Tricks of the trade.

Mathematica Journal, 6(3):22– 27, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i3/tutorials/tricks/
contents/63tricks.nb;   http:
//www.mathematica-journal.
com/issue/v6i3/tutorials/tricks/
contents/63tricks.pdf;   http:
//www.mathematica-journal.
com/issue/v6i3/tutorials/tricks/
index.html.

Abbott:1996:TTd

[Abb96d] Paul Abbott. Tricks of the trade. Mathematica Journal, 6 (4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i4/tutorials/tricks/
index.html.

Abbott:1996:IOa

[Abb96e] Paul Abbott. In[] and Out[].

Mathematica Journal, 6(1):??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i1/tutorials/inout/
index.html.

Abbott:1996:IOb

[Abb96f] Paul Abbott. In[] and Out[]. Mathematica Journal, 6(2):??, Spring 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i2/tutorials/inout/
index.html;   http://www.mathematica-
journal.com/issue/v6i2/tutorials/
inout/inout62.nb.

Abbott:1996:IOc

[Abb96g] Paul Abbott. In[] and Out[]. Mathematica Journal, 6(3):14– 21, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i3/tutorials/inout/
contents/63inandout.nb;   http:
//www.mathematica-journal.
com/issue/v6i3/tutorials/inout/
contents/63inout.pdf;   http:
//www.mathematica-journal.
com/issue/v6i3/tutorials/inout/
index.html.

Abbott:1996:IOd

[Abb96h] Paul Abbott. In[] and Out[]. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i4/tutorials/inout/
index.html.

Abbott:1997:TT

[Abb97a] Paul Abbott. Tricks of the trade. Mathematica Journal, 7(1): ??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v7i1/tutorials/tricks/
index.html.

Abbott:1997:TTC

[Abb97b] Paul Abbott. Tricks of the trade: Computing complex conjugates, and other programming techniques. Mathematica Journal, 7(3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1997:TTF

[Abb97c] Paul Abbott. Tricks of the trade:

Formatted tables and statistical distributions. Mathematica Journal, 7(4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1997:IO

[Abb97d] Paul Abbott. In[] and Out[].

Mathematica Journal, 7(1):??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v7i1/tutorials/inout/
index.html.

Abbott:1997:EAYa

[Abb97e] Paul Abbott. In[] and Out[]:

Experts answer your questions. Mathematica Journal, 7(3):??, ???? 1997. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1997:EAYb

[Abb97f] Paul Abbott. In[] and Out[]: Experts answer your questions. Mathematica Journal, 7(4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1998:TT

[Abb98a] Paul Abbott. Tricks of the trade. Mathematica Journal, 8 (1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1998:TTN

[Abb98b] Paul Abbott. Tricks of the trade: Numerical differentiation using Fourier, Files palette, smooth interpolation, and more. Mathematica Journal, 8(3):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1998:EAYa

[Abb98c] Paul Abbott. In[] and Out[]: Experts answer your questions. Mathematica Journal, 8(1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1998:EAYb

[Abb98d] Paul Abbott. In[] and Out[]: Experts answer your questions. Mathematica Journal, 8(3):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1998:EAYc

[Abb98e] Paul Abbott. In[] and Out[]:

Experts answer your questions. Mathematica Journal, 8(4):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1999:TTI

[Abb99a] Paul Abbott. Tricks of the trade: Interpolation with noise, using reduce to solve the Kuhn– Tucker equations, Legendre– Gauss quadrature, Pascal matrices, and generalized Padé approximation. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1999:TTN

[Abb99b] Paul Abbott. Tricks of the trade:

Nondecreasing function, proofs of inequalities, and explicit roots of transcendental equations. Mathematica Journal, 9(3):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1999:TTS

[Abb99c] Paul Abbott. Tricks of the trade: Symbolic analysis of DNA sequences, generating barcodes, computing your tax, and ParentForm. Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1999:EAYa

[Abb99d] Paul Abbott. In[] and Out[]:

Experts answer your questions.

Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1999:EAYb

[Abb99e] Paul Abbott. In[] and Out[]: Experts answer your questions. Mathematica Journal, 9(3):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:1999:EAYc

[Abb99f] Paul Abbott. In[] and Out[]: Experts answer your questions. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:2005:ISI

[Abb05a] Paul Abbott. Introduction to the Seventhg International Mathematica Symposium, (Perth, Western Australia, August 5–8, 2005). Mathematica Journal, 10 (2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:2005:TTS

[Abb05b] Paul Abbott. Tricks of the trade: Sum-free set, Google search for Mathematica, asymptotic expansion and π Hadamard regularization. Mathematica Journal, 10 (3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:2005:EAY

[Abb05c] Paul Abbott. In[] and Out[]:

Experts answer your questions. Mathematica Journal, 10(2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Abbott:2006:PE

[Abb06] Paul Abbott. On the perimeter of an ellipse. Mathematica Journal, 11(2):172–185, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v11i2/contents/Abbott/
Abbott.nb;   http://www.mathematica-
journal.com/issue/v11i2/contents/
Abbott/Abbott.pdf.

Agnew:2010:TCS

[ABBS10] Alfonso F. Agnew, Alexandru Bobe, Wladimir G. Boskoff, and Bogdan D. Suceava. Tzitzeica curves and surfaces. Mathematica Journal, 12(1):3:1–3:18, ???? 2010. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/data/uploads/2010/10/Agnew.
pdf.

Arwini:1999:NIR

[AD99] Khadiga Arwini and C. T. J. Dodson. Neighbourhoods of independence for random processes via information geometry. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Adamchik:1992:LCF

[Ada92a] V. Adamchik. Limits of continued fractions and nested radicals. Mathematica Journal, 2(2):54–57, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Adamchik:1991:LCF

[Ada92b] Victor S. Adamchik. Limits of continued fractions and nested radicals. Mathematica Journal, 2(2):54–57, ???? 1992. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v2i2/article/
                                        adamchik/index.html.

Adamchik:1996:DI

[Ada96] Victor Adamchik. Definite integration. Mathematica Journal, 6 (4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i4/version3.0/
                                        adamchik/index.html.

Akers:1998:B

[AG98] David L. Akers and Robert N. Goldberg. BioEqCalc. Mathematica Journal, 8(1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Aguirregabiria:1990:CRE

[Agu90] Juan M. Aguirregabiria. Computing the Ricci and Einstein tensors. Mathematica Journal, 1(2): 51–54, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/article/juanma/
index.html.

Allen:1991:SQM

[AH92] Arnold O. Allen and Gary Hynes. Solving a queueing model with Mathematica. Mathematica Journal, 1(3):108–112, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i3/article/allen/
index.html.

Aizenberg:1997:ARH

[Aiz97]  Lev Aizenberg.  Approach-

ing the Riemann Hypothesis with Mathematica. Mathematica Journal, 7(1):??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v7i1/article/aizenberg/
index.html.

Abbott:1991:IOa

[AL91] Paul C. Abbott and Silvio Levy. In[] and Out[]. Mathematica Journal, 1(3):??, Winter 1991. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i3/tutorials/inout/
index.html.

Abbott:1991:OLa

[AL92a] Paul C. Abbott and Silvio Levy. One-liners. Mathematica Journal, 1(3):??, Winter 1992. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i3/tutorials/oneliners/
                                        index.html.

Abbott:1991:OLb

[AL92b] Paul C. Abbott and Silvio Levy. One-liners. Mathematica Journal, 2(1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/tutorials/oneliners/
                                        index.html.

Abbott:1991:TTa

[AL92c] Paul C. Abbott and Silvio Levy. Tricks of the trade. Mathematica Journal, 2(1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/tutorials/tricks/
                                        index.html.

Allen:2013:DDG

[All13]  Todd D. Allen.  Detect-

ing differential gene expression using Affymetrix microarrays. Mathematica Journal, 15(??):??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/2013/11/detecting-differential-
                                        gene-expression-using-affymetrix-
                                        microarrays/.

Antchev:1997:IEM

[AMP97] Kostadin Antchev, Jari Multisilta, and Seppo Pohjolainen. Interactive exercises on matrix algebra. Mathematica Journal, 7 (3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Andres:1990:NNW

[And90]  Clay Andres.  Neural net

work challenges Mathematica. Mathematica Journal, 1(2):12– 15, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/tutorials/inout/
index.html.

Anonymous:1991:MGC

[Ano92] Anonymous. 1992 Mathematica graphics competition. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1994:CEa

[Ano94a] Anonymous. Calendar of events.

Mathematica Journal, 4(1):??, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1994:CEb

[Ano94b] Anonymous. Calendar of events.

Mathematica Journal, 4(2):??, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1994:MNa

[Ano94c] Anonymous. Mathematica news.

Mathematica Journal, 4(1):??, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1994:MNb

[Ano94d] Anonymous. Mathematica news. Mathematica Journal, 4(2):??, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1994:PAa

[Ano94e] Anonymous. Program announcements. Mathematica Journal, 4 (1):??, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1994:PAb

[Ano94f] Anonymous. Program announcements. Mathematica Journal, 4 (2):??, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1995:CE

[Ano95a] Anonymous. Calendar of events. Mathematica Journal, 5(4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1995:IV

[Ano95b] Anonymous. Index to volume 5. Mathematica Journal, 5(4): ??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1995:MN

[Ano95c] Anonymous. Mathematica news. Mathematica Journal, 5(4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1997:DI

[Ano97a] Anonymous. The area of a random triangle. Mathematica Journal, 7(2):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1997:D

[Ano97b] Anonymous. The design of the Mathematica fonts. Mathematica Journal, 7(2):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1997:H

[Ano97c] Anonymous. Version 3.0 in the classroom. Mathematica Journal, 7(2):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1997:FG

[Ano97d] Anonymous. Visualization of complex functions. Mathematica Journal, 7(2):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:1998:TTI

[Ano98]  Anonymous.  Tricks of the

trade: Inflight conundrums, integral equations, matrix differential operator. Mathematica Journal, 8(4):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:2006:FPA

[Ano06] Anonymous. Free and paperless — all issues of The Mathematica Journal are now available in an online-only format with free, unlimited access for everyone. Mathematica Journal, 11 (3):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Adickes:1991:STU

[ARAM92] M. D. Adickes, R. H. Rucker, M. R. Anderson, and W. C. Moor. Structuring tutorials using Mathematica: Educational theory and practice. Mathematica Journal, 1(3):86–91, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i3/article/adickes/
index.html.

Abad:1995:CRC

[AS95] Julio Abad and Javier Sesma. Computation of the regular confluent hypergeometric function. Mathematica Journal, 5 (4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/article/abad/
index.html.

Auluck:2012:IPT

[Aul12] S. K. H. Auluck. On the integral of the product of three Bessel functions over an infinite domain: Fourier-space representation of nonlinear dynamics of continuous media in cylindrical geometry. Mathematica Journal, 14(??):??, ???? 2012. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/2012/12/on-the-
integral-of-the-product-of-
three-bessel-functions-over-
an-infinite-domain/.

Akritas:2006:SAF

[AUV06] Alkiviadis G. Akritas, Jerry Uhl, and Panagiotis S. Vigklas. On some applications of the Fast Discrete Fourier Transform. Mathematica Journal, 11(1):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Bagis:2013:GMC

[Bag13] N. D. Bagis. A general method for constructing Ramanujan-type formulas for powers of 1/π. Mathematica Journal, 15(??):??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2013/08/a-general-method-
for-constructing-ramanujan-
type-formulas-for-powers-of-
1-over-pi/.

Balkin:1994:EFT

[Bal94]  Sandy D. Balkin.  Educa-

tion forum: Taking calculus with Mathematica. Mathematica Journal, 4(2):52–53, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i2/columns/forum/
52-53_balkin42.mj.pdf;   http:
//www.mathematica-journal.

                                        com/issue/v4i2/columns/forum/
                                        index.html.

Barrere:2006:AAM

[Bar06] Rémi Barrère. An algorithmic approach to manifolds. Mathematica Journal, 11(2):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v11i2/Barrere.html.

Bulmer:1996:IPM

[BC96] Michael Bulmer and Michael Carter. Integer programming with Mathematica. Mathematica Journal, 6(3):28–36, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i3/tutorials/bulmer/
                                        contents/63bulmer.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v6i3/tutorials/bulmer/
                                        index.html.

Businger:1994:CMS

[BCDH94] W. Businger, P.-A. Chevalier, N. Droux, and W. Hett. Computing minimal surfaces on a transputer network. Mathematica Journal, 4(2):70–75, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i2/article/businger/
                                        index.html.

Beck:1993:GM

[Bec93] George Beck. Geometry in motion. Mathematica Journal, 3 (4):??, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Belsley:1990:EMA

[Bel90a] D. Belsley. Econometrics.m: a basic package for econometrics. Mathematica Journal, 1(1):95– 101, Summer 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Belsley:1990:EMB

[Bel90b] David A. Belsley. Econometrics.m: a basic package for econometrics. Mathematica Journal, 1(1):95–101, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i1/features/belsley/
index.html.

Belknap:1991:CKS

[Bel92] Steven Belknap. The Chicago Kinetic Simulator. Mathematica Journal, 1(4):68–86, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i4/article/belknap/
index.html.

Benayat:2005:GRS

[Ben05] Djilali Benayat. Graphing on the Riemann sphere. Mathematica Journal, 10(4):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Blachman:1992:LMI

[Bla92a] Nancy Blachman. Learning Mathematica: Importing and exporting data. Mathematica Journal, 2(3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i3/tutorials/blachman/
index.html.

Blachman:1992:LMT

[Bla92b] Nancy Blachman. Learning Mathematica: Test your knowledge. Mathematica Journal, 2 (4):??, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i4/tutorials/blachman/
index.html.

Blachman:1991:NFL

[Bla92c] Nancy Blachman. Nonlinear fitting, looping and recursion. Mathematica Journal, 2(2):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Blachman:1993:LMM

[Bla93a] Nancy Blachman. Learning Mathematica: More on importing data. Mathematica Journal, 3(1): ??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i1/columns/blachman/
index.html.

Blachman:1993:LMT

[Bla93b] Nancy Blachman. Learning Mathematica: Two-dimensional Mathematica graphics. Mathematica Journal, 3(3):??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/tutorials/blachman/
index.html.

Blachman:1993:LMU

[Bla93c] Nancy Blachman. Learning Mathematica: Using the standard packages. Mathematica Journal, 3(2):31–35, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i2/tutorials/blachman/
index.html.

Blachman:1993:TNR

[Bla93d] Nelson Blachman. There are nine regular solids, not just five. Mathematica Journal, 3(2): ??, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i2/article/blachman/
index.html.

Blachman:1998:DR

[Bla98] Nancy Blachman. Demystifying rules. Mathematica Journal, 8(4): ??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.

com/issue/v8i4/features/blachman/
index.html.

Bakshee:1994:RSC

[BM94] Igor S. Bakshee and Toshimitsu Musha. Random series in computer graphics. Mathematica Journal, 4(1):9–10, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i1/graphics/09-
10_galleryText.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i1/graphics/index.
html.

Baumann:1999:EOO

[BM99] Gerd Baumann and Michal Mnuk. Elements — an objectoriented approach to industrial software development. Mathematica Journal, 10(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Bar-Natan:1991:RDS

[BN92] Dror Bar-Natan. Random-dot stereograms. Mathematica Journal, 1(3):69–75, Winter 1992. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Bocharov:1993:SSN

[Boc93] Alexei V. Bocharov. Symbolic solvers for nonlinear differential equations. Mathematica Journal, 3(2):??, Spring 1993. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v3i2/article/
bocharov/index.html.

Bonham:2012:SII

[Bon12] Russell A. Bonham. Some integrals involving symmetric-top eigenfunctions. Mathematica Journal, 14(??):??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/2012/05/some-integrals-
involving-symmetric-top-eigenfunctions/

.

Borkovsky:1991:M

[Bor92] Arkady Borkovsky. MathLink.

Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Brown:1990:CMC

[BPU90] Don Brown, Horacio Porta, and Jerry Uhl. Calculus & Mathematica: Courseware for the nineties. Mathematica Journal, 1(1):43– 50, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i1/features/brown/
index.html.

Bringslid:1999:XC

[Bri99] Odd Bringslid. The Xmath Calculator. Mathematica Journal, 10(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Benninga:1993:INO

[BSS93] Simon Benninga, Raz Steinmetz, and John Stroughair. Implementing numerical option pricing models. Mathematica Journal, 3(4):

66–73, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i4/article/benninga/
index.html.

Berdnikov:1996:NMP

[BTC96] A. S. Berdnikov, S. B. Trutia, and A. Compagner. Notebook: A MathLink program for high-quality random numbers. Mathematica Journal, 6(3):65– 69, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i3/article/berdnikov/
contents/63berdnikov.nb;
http://www.mathematica-journal.
com/issue/v6i3/article/berdnikov/
contents/63berdnikov.pdf;
http://www.mathematica-journal.
com/issue/v6i3/article/berdnikov/
index.html.

Buchberger:1998:TPS

[Buc98] Bruno Buchberger. Theorema: a proving system based on Mathematica. Mathematica Journal, 8(2):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Carter:1994:LPM

[Car94] Michael Carter. Linear programming with Mathematica: The simplex algorithm. Mathematica Journal, 4(4):??, Fall 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v4i4/tutorials/
carter/index.html.

Carter:1995:LPM

[Car95] Michael Carter. Linear programming with Mathematica: Sensitivity analysis. Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i1/tutorials/carter/
index.html.

Carter:2011:TQS

[Car11]  Michael Carter.  A tool-

box for quasirandom simulation. Mathematica Journal, 13(??):??, ???? 2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2011/12/a-toolbox-for-
quasirandom-simulation/.

Cassel:2014:PPS

[Cas14]  John Cassel.  Probabilistic

programming with stochastic memoization: Exploring fractal curves implementing nonparametric Bayesian inference. Mathematica Journal, 16(??):??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2014/01/probabilistic-
programming-with-stochastic-
memoization/.

Chartier:2011:CMD

[CC11] Sylvain Chartier and Denis Cousineau. Computing mixeddesign (split-plot) ANOVA. Mathematica Journal, 13(??):??, ????

2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (elec-

                                        tronic).  URL http://www.
                                        mathematica-journal.com/2011/
                                        10/computing-mixed-design-
                                        split-plot-anova/.

Cohen:1992:DAL

[CD92a] J. Cohen and D. DeBaun. Discrete approximations of linear functionals. Mathematica Journal, 2(2):62–65, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Cohen:1991:DAL

[CD92b] Jack K. Cohen and David R. DeBaun. Discrete approximation of linear functions. Mathematica Journal, 2(2):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i2/article/cohen/
                                        index.html.

Castillo:1995:SFE

[CGI95] Enrique Castillo, José Manuel Gutiérrez, and Andrés Iglesias. Solving a functional equation. Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i1/article/castillo/
                                        index.html.

Czaja:2011:AWE

[CGW11] Wojciech Czaja, Zdzislaw A. Golda, and Andrzej Woszczyna. The acoustic wave equation in the expanding universe: Sachs– Wolfe Theorem. Mathematica Journal, 13(??):??, ???? 2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/2011/10/the-acoustic-
wave-equation-in-the-expanding-
universe-sachs-wolfe-theorem/

.

Chapnick:1994:AIG

[Cha94] Philip Chapnick. Book review:

Ad Infinitum: The Ghost in Turing’s Machine and Zero to Lazy Eight: Playing with infinity. Mathematica Journal, 4(1):20– 21, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i1/reviews/chapnick/
20-21_Chapnick.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i1/reviews/chapnick/
index.html.

Chandler:2005:NSS

[Cha05] Seth J. Chandler. The network structure of Supreme Court jurisprudence. Mathematica Journal, 10(3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Christensen:1992:MBC

[Chr92a] Steven M. Christensen. From the E-mail bag: Computing determinants. Mathematica Journal, 2 (4):??, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.

                                        com/issue/v2i4/tutorials/christensen/
                                        index.html.

Christensen:1991:MBT

[Chr92b] Steven M. Christensen. From the E-mail bag: Topics in PostScript and electronic mail. Mathematica Journal, 2(2):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i2/tutorials/christensen/
                                        index.html.

Christensen:1993:MBF

[Chr93a] Steven M. Christensen. From the E-mail bag: Frequently asked questions. Mathematica Journal, 3(1):??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v3i1/columns/christensen/
                                        index.html.

Christensen:1993:MBH

[Chr93b] Steven M. Christensen. From the E-mail bag: History, nohup, and the end.m. Mathematica Journal, 3(2):??, Spring 1993. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v3i2/columns/
                                        christensen/index.html.

Chandler:1998:BLU

[CJ98] Seth J. Chandler and Christian Jacob. Behavior and learning under law. Mathematica Journal, 8(2):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Cohen:1992:PLS

[Coh92a] J. Cohen. Packages for logic and set theory. Mathematica Journal, 2(1):91–93, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Cohen:1991:PLS

[Coh92b] Jack K. Cohen. Packages for logic and set theory. Mathematica Journal, 2(1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v2i1/article/
cohen/index.html.

Cohen:1995:CND

[Coh95a] Jack K. Cohen. Classroom notes:

Disordered lattices and methods for calculating pi. Mathematica Journal, 5(3):??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i3/columns/cohen/
index.html.

Cohen:1995:CNR

[Coh95b] Jack K. Cohen. Classroom notes: Rational-length vectors and random matrices. Mathematica Journal, 5(2):??, Spring 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i2/columns/cohen/
index.html.

Cohen:1995:CNS

[Coh95c] Jack K. Cohen. Classrooom notes: Symbolic integration and asymptotic expansions. Mathematica Journal, 5(4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (elec-

                                        tronic).  URL http://www.
                                        mathematica-journal.com/issue/
                                        v5i4/columns/cohen/index.html.

Cohen:1996:CNF

[Coh96]  Jack K. Cohen.  Classroom

notes: Fluid flow past a cylinder. Mathematica Journal, 6(1): ??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i1/columns/cohen/
                                        index.html.

Cook:1991:S

[Coo92] Matthew Cook. Sound. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Coppola:1991:WS

[Cop92] Vincent T. Coppola. Working with subexpressions. Mathematica Journal, 1(4):??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v1i4/tutorials/
                                        coppola/index.html.

Courant:1993:MGT

[Cou93]  Ted Courant.  Mathemat-

ica games: Towers of power: a potent paradox. Mathematica Journal, 3(3):??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/columns/courant/
index.html.

Courant:1995:PGS

[Cou95]  Ted Courant.  Pictures of

geometric series. Mathematica Journal, 5(3):??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i3/graphics/courant/
index.html.

Cousineau:2013:CRB

[Cou13] Denis Cousineau. Controlling robots built with the LEGO(R) MINDSTORMS(R) NXT brick. Mathematica Journal, 15(??):??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2013/03/controlling-robots-
built-with-the-lego-mindstorms-
nxt-brick/.

Cowen:2013:UBC

[Cow13] Robert Cowen. Using Boolean computation to solve some problems from Ramsey theory. Mathematica Journal, 15(??):??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2013/10/using-boolean-
computation-to-solve-some-

                                        problems-from-ramsey-theory/

.

Christensen:1990:MSP

[CP90] Steven M. Christensen and Leonard Parker. MathTensor: a system for performing tensor analysis by computer. Mathematica Journal, 1(1):51– 61, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i1/features/christensen/
                                        index.html.

Cheng:1994:AIL

[CSA94] Alexander H-D. Cheng, Paston Sidauruk, and Younane Abousleiman. Approximate inversion of the Laplace transform. Mathematica Journal, 4(2):76– 82, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i2/article/cheng/
                                        index.html.

Cabrera:1992:ISM

[CW92a] J. Cabrera and A. Wilks. An interface from S to Mathematica. Mathematica Journal, 2(1):66–74, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Cabrera:1991:IM

[CW92b] Javier F. Cabrera and Alan R. Wilks. An interface from S to Mathematica. Mathematica Journal, 2(1):??, ???? 1992. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v2i1/article/
cabrera/index.html.

Clarke:1993:ATP

[CZ93] Edmund Clarke and Xudong Zhao. Analytica: a theorem prover for Mathematica. Mathematica Journal, 3(1):??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i1/article/clarke/
index.html.

Dai:1998:FEM

[Dai98] Naci M. Dai. Finite element method. Mathematica Journal, 8(1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

DeSmedt:1999:AA

[De 99] Stany De Smedt. p-adic arithmetic. Mathematica Journal, 9 (2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Dudas:1991:AQM

[DEH92] Magdalena M. Dudas, Walter C.

Ermler, and Hsiuchin C. Hsieh. Application of quantum mechanical perturbation theory to molecular vibrational-rotational analysis. Mathematica Journal, 2(2): 66–69, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i2/article/dudas/
index.html.

Dudas:1992:AQM

[DHE92] M. Dudas, H. Hsieh, and W. Ermler. Application of quantum mechanical perturbation theory to molecular vibrationalrotational analysis. Mathematica Journal, 2(2):66–69, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Dickson:1990:MS

[Dic90] Stewart Dickson. Minimal surfaces. Mathematica Journal, 1 (1):??, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Dickson:1991:MHS

[Dic92] Stewart Dickson. Many-handled surfaces. Mathematica Journal, 1 (4):??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i4/graphics/index.
                                        html.

Dill:1991:CPT

[Dil92a] D. Dill. Creating packages with TEX/Mathematica. Mathematica Journal, 1(4):16–17, Spring 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Dill:1991:ITM

[Dil92b] Dan Dill. Interactive TEX/ Mathematica documents. Mathematica Journal, 1(3):104–107, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.

com/issue/v1i3/article/dill/
index.html.

Dickhaut:1991:PFN

[DK92a] John Dickhaut and Todd Kaplan.

A program for finding Nash equilibria. Mathematica Journal, 1(4): ??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i4/article/dickhaut/
index.html.

Dikhaut:1991:PFN

[DK92b] J. Dikhaut and T. Kaplan. A program for finding Nash equilibria. Mathematica Journal, 1(4):87–93, Spring 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

deOliveira:2010:DMS

[dO10] Pedro P. B. de Oliveira. DRIMA:

a minimal system for probing the dynamics of change in a reactive multi-agent setting. Mathematica Journal, 12(1):1:1–1:18, ???? 2010. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/data/uploads/2010/10/deOliveira.
pdf.

deOliveira:2014:RFC

[dOV14] Pedro P. B. de Oliveira and Maurício Verardo. Representing families of cellular automata rules. Mathematica Journal, 16 (??):??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.

                                        com/2014/08/representing-families-
                                        of-cellular-automata-rules/.

Diggle:2005:IBM

[DSM05] Art Diggle, Moin Salam, and Marta Monjardino. Individualbased models of the spread of disease, weeds, and insects. Mathematica Journal, 10(2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Dasgupta:1998:MFF

[DT98] Gautam Dasgupta and Jacques Treil. Maxillo-facial frames: Finite element shapes. Mathematica Journal, 8(2):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Dubisch:1990:TKN

[Dub90a] Russell J. Dubisch. The tool kit: a notebook subclass. Mathematica Journal, 1 (2):??, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i2/article/dubisch/
                                        index.html.

Dubish:1990:TKA

[Dub90b] R. Dubish. The tool kit: a notebook subclass. Mathematica Journal, 1(2):55–64, Fall 1990. ISSN 1047-5974 (print), 1097-1610 (electronic).

Ebneter:1990:G

[Ebn90] Kate Ebneter. Geode. Mathematica Journal, 1(1):??, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Edelen:1991:ECV

[Ede92a] Dominic G. B. Edelen. Exterior calculus for vectors and forms. Mathematica Journal, 1(3):92– 95, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i3/article/edelen/
index.html.

Edelman:1992:CPC

[Ede92b] Alan Edelman. The complete pivoting conjecture for Gaussian elimination is false. Mathematica Journal, 2(2):58–61, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Edelman:1991:CPC

[Ede92c] Alan Edelman. The complete pivoting conjecture for Gaussian elimination is false. Mathematica Journal, 2(2):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i2/article/edelman/
index.html.

Ediger:1990:GNM

[Edi90] M. N. Ediger. A Gauss–Newton method for nonlinear regression. Mathematica Journal, 1(2):42– 44, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.

                                        com/issue/v1i2/article/ediger/
                                        index.html.

Edwards:1993:TT

[Edw93] C. Henry Edwards. Twisted tubes. Mathematica Journal, 3(1):??, Winter 1993. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v3i1/graphics/
                                        index.html.

Edwards:1994:LML

[Edw94] C. Henry Edwards. Ladders, moats, and Lagrange multipliers. Mathematica Journal, 4(1): ??, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i1/tutorials/edwards/
                                        48-52_Edwards.mj.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v4i1/tutorials/edwards/
                                        index.html.

Edwards:1997:NNC

[Edw97] C. Henry Edwards. Newton’s nose-cone problem. Mathematica Journal, 7(1):??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v7i1/article/edwards/
                                        index.html.

Evans:1990:STA

[EMM90] Brian L. Evans, James H. Mc-Clellan, and Wallace B. McClure. Symbolic transforms with applications to signal processing.

Mathematica Journal, 1(2):70– 80, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/article/evans/
index.html.

Enderlein:1997:PDC

[End97] Jörg Enderlein. A package for displaying crystal morphology. Mathematica Journal, 7(1):??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v7i1/article/enderlein/
index.html.

Entacher:1999:BUD

[Ent99] Karl Entacher. On the beauty of uniform distribution modulo one. Mathematica Journal, 9 (3):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Eisenberg:2006:VCF

[EP06] Murray Eisenberg and David J. M. Park, Jr. Visualizing complex functions with the presentations application. Mathematica Journal, 11(2):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v11i2/EisenbergPark.
html.

Erdos:2006:SRM

[Erd06] Gábor Erdős. Substitutions and replacements in mechanism prototyping. Mathematica Journal, 11(2):??, ???? 2006. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v11i2/Erdos.
                                        html.

Falloon:2005:MMM

[Fal05] Peter Falloon. A mode-matching method for multichannel scattering problems. Mathematica Journal, 10(3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Farley:1991:FUA

[Far92] Anthony R. Farley. The fractal universe of Andrei Linde. Mathematica Journal, 2(1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/graphics/farley/
                                        index.html.

Fearn:2005:AAU

[Fea05] Richard L. Fearn. Airfoil aerodynamics using panel methods. Mathematica Journal, 10(4):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Fritzson:2005:MSC

[FES05] Peter Fritzson, Vadim Engelson, and Krishnamurthy Sheshadri. MathCode: a system for C++ or Fortran code generation from Mathematica. Mathematica Journal, 10(4):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Fielder:1999:GRP

[Fie99] Daniel C. Fielder. A generator of rook polynomials. Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Foale:1997:MM

[Foa97] Michael Foale. Mir and Mathematica. Mathematica Journal, 7(3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Fairchild:1995:PSA

[FPP95] Steve Fairchild, Ruth Pachter, and Ronald Perrin. Protein structure analysis and prediction. Mathematica Journal, 5 (4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/article/fairchild/
index.html.

Freeman:1993:AIN

[Fre93a] James Freeman. Artificial intelligence: a neural network solution to the n-queens problem. Mathematica Journal, 3(3):??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/columns/freeman/
index.html.

Freeman:1993:AIS

[Fre93b] James Freeman. Artificial intelligence: Simulating a basic genetic algorithm. Mathematica Journal, 3(2):??, Spring 1993. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v3i2/columns/
freeman/index.html.

Freeman:1994:AIF

[Fre94]  James A. Freeman.  Artifi-

cial intelligence: Fuzzy systems for control applications: The truck backer-upper. Mathematica Journal, 4(1):64–69, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i1/columns/freeman/
64-69_Jim41.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i1/columns/freeman/
index.html.

Fritzson:1999:MOO

[Fri99] Peter Fritzson. MathModelica – 0 an object-oriented mathematical modeling and simulation environment. Mathematica Journal, 10(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Fultz:1994:TSC

[Ful94a] John Fultz. Tech support: Curve fits, MathLink for Excel, and color. Mathematica Journal, 4 (4):??, Fall 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i4/columns/techsupport/
index.html.

Fultz:1994:TSG

[Ful94b] John Fultz. Tech support: Graphics labels, ReadList, and imports and exports. Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i3/columns/techsupport/
index.html.

Fultz:1995:TSL

[Ful95a] John Fultz. Tech support: Legends and MathLink questions. Mathematica Journal, 5(3):??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i3/columns/techsupport/
index.html.

Fultz:1995:TSN

[Ful95b] John Fultz. Tech support: Numerical precision and tick marks. Mathematica Journal, 5(2):??, Spring 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i2/columns/techsupport/
index.html.

Fultz:1995:TSP

[Ful95c]  John Fultz.  Tech support:

Package programming. Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.

com/issue/v5i1/columns/techsupport/
index.html.

Fultz:1996:TSA

[Ful96a] John Fultz. Tech support: Arrows and local variables. Mathematica Journal, 6(2):52–59, Spring 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i2/columns/techsupport/
index.html;   http://www.mathematica-
journal.com/issue/v6i2/columns/
techsupport/tech62.pdf.

Fultz:1996:TSV

[Ful96b] John Fultz. Tech support: Visualization of 3D data and numerical integration. Mathematica Journal, 6(1):??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i1/columns/techsupport/
index.html.

Fine:1991:SPS

[FW92] Jerry Fine and Lee R. Waite. Spatial problem solving in an integrated first-year curriculum. Mathematica Journal, 1(4):62– 67, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i4/article/fine/
index.html.

Gayley:1993:MTC

[Gay93a] Todd Gayley. The Mathematica toolbox: CleanSlate: a package for clearing symbols and freeing memory. Mathematica Journal, 3(3):??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/columns/toolbox/
index.html.

Gayley:1993:MTE

[Gay93b] Todd Gayley. The Mathematica toolbox: Enhanced error messages. Mathematica Journal, 3 (4):??, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i4/columns/toolbox/
index.html.

Gayley:1993:TSX

[Gay93c] Todd Gayley. Tech support: The X front end and UNIX installation. Mathematica Journal, 3 (4):??, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i4/columns/techsupport/
index.html;   http://www.mathematica-
journal.com/issue/v4i2/columns/
techsupport/54-61_gayley42.
mj.pdf.

Gayley:1994:MTM

[Gay94a] Todd Gayley. The Mathematica toolbox: a MathLink program for accessing binary files. Mathematica Journal, 4(2):44– 51, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i2/columns/toolbox/

                                        44-51_gayley42.mj.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v4i2/columns/toolbox/
                                        index.html.

Gayley:1994:TSC

[Gay94b] Todd Gayley. Tech support: Customizing the X front end, handling large lists, and generating surface plots. Mathematica Journal, 4(2):54–61, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i2/columns/techsupport/
                                        index.html.

Gayley:1994:TSR

[Gay94c] Todd Gayley. Tech support: Remote kernels, ComplexExpand, and memory use under Windows. Mathematica Journal, 4(1):70– 73, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i1/columns/techsupport/
                                        70-73_Todd41.mj.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v4i1/columns/techsupport/
                                        index.html.

Gayley:1995:MTM

[Gay95] Todd Gayley. The Mathematica toolbox: a Mathematica profiler. Mathematica Journal, 5(3): ??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i3/columns/toolbox/
                                        index.html.

Gayley:1999:BUI

[Gay99] Todd Gayley. Building user interfaces using J/Link. Mathematica Journal, 9(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Gehrig:1997:O

[Geh97] David Gehrig. The Orloj. Mathematica Journal, 7(4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Germundsson:1997:MV

[Ger97] Roger Germundsson. Mathematica Version 4. Mathematica Journal, 7(4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Gray:1997:IMM

[GG97] Theodore Gray and Jerry Glynn.

Image manipulation with Mathematica 4. Mathematica Journal, 7(4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Gutierrez:1997:GRF

[GIRR97] José Manuel Gutiérrez, Andrés Iglesias, Miguel Angel Rodríguez, and Victor José Rodríguez. Generating and rendering fractal images. Mathematica Journal, 7(1): ??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v7i1/graphics/index.
html.

Goffinet:1995:WMT

[Gof95]  Daniel Goffinet.  A wooden

Möbius trefoil knot. Mathematica Journal, 5(4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v5i4/article/
                                        goffinet/index.html.

Gray:1990:SFS

[Gra90] Theodore W. Gray. The seven faces of sin. Mathematica Journal, 1(1):??, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Gray:1991:MFE

[Gra92] Theo Gray. Macintosh front end. Mathematica Journal, 1 (3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Gray:1996:FE

[Gra96] Theo Gray. The front end. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i4/version3.0/
                                        gray/index.html.

Gregg:1998:LID

[Gre98] Joe Gregg. Leibniz: Implementing a drag-and-drop calculator. Mathematica Journal, 8 (3):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Gunn:1990:CMS

[Gun90] Charlie Gunn. Computers in mathematics: snapshots of a possible future. Mathematica Journal, 1(2):24–25, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/tutorials/inout/
index.html.

Garnerin:1991:VTR

[GVV92] Philippe Garnerin, Alain-Jacques Valleron, and Jacques Vidal. A visualization tool for representing epidemic spread. Mathematica Journal, 1(4):59–61, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i4/article/garnerin/
index.html.

Harris:1994:REP

[Har94a] Jason F. Harris. Rearranging expressions by patterns. Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i3/article/harris/
index.html.

Hart:1994:SRF

[Har94b] David Hart. Software review:

Function browsers and help stacks. Mathematica Journal, 4 (1):26–27, Winter 1994. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v4i1/reviews/
                                        hart/26-27_hart.mj.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v4i1/reviews/hart/
                                        index.html.

Harris:1997:SSP

[Har97a] Jason Harris. Semantica: Semantic pattern matching in Mathematica. Mathematica Journal, 7(3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

hart:1997:Z

[Har97b] George W. Hart. Zonohedrification. Mathematica Journal, 7 (3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Hayes:1992:PPR

[Hay92a] A. Hayes. Performance of programs revisited. Mathematica Journal, 2(2):50–53, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Hayes:1991:PPR

[Hay92b] Allan Hayes. Performance of programs revisited. Mathematica Journal, 2(2):50–53, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i2/article/hayes/
                                        index.html.

Hayes:1995:EEP

[Hay95] Allan Hayes. Experiments in efficient programming. Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i1/tutorials/hayes/
index.html.

Hebert:1990:D

[Héb90]  Jean-Pierre Hébert.  Draw-

ings. Mathematica Journal, 1 (2):??, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Hebert:1997:MDS

[Héb97] Jean-Pierre Hébert. Mathematicadesigned sand tracings. Mathematica Journal, 7(3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Helzer:1992:CSE

[Hel92]  Garry Helzer.  Constructions

with straight edge and compass. Mathematica Journal, 2(3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i3/article/helzer/
index.html.

Helzer:1995:GB

[Hel95] Garry Helzer. Gröbner bases.

Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i1/article/helzer/
index.html.

Helzer:1998:CP

[Hel98] Garry Helzer. Clipping polygons. Mathematica Journal, 8 (4):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Hertel:1998:QTM

[Her98] Joachim Hertel. Quantum Turing Machine simulator. Mathematica Journal, 8(3):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Herod:1999:LAD

[Her99] Scott A. Herod. Learning about differential equations from their symmetries. Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Hertel:2006:CUR

[Her06] Joachim Hertel. Computing the uncomputable Rado sigma function. Mathematica Journal, 11 (2):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v11i2/Hertel.html.

Horn:1994:AGD

[HL94] Geir Horn and Svein Linge. Analytical generation of the dynamical equations for mechanical manipulators. Mathematica Journal, 4(4):??, Fall 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i4/article/horn/
                                        index.html.

Hector:1996:AES

[HLF96] Louis G. Hector, Jr., Kenneth B.

Lippert, and Joseph M. Fridy. Analysis of engineering surfaces. Mathematica Journal, 6(2):72– 79, Spring 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i2/article/hector/
hector.pdf;   http://www.mathematica-
journal.com/issue/v6i2/article/
hector/index.html.

Hector:1996:CT

[HN96] Louis G. Hector, Jr. and Kenneth B. Lippert Notebook. Coin topographies. Mathematica Journal, 6(3):10–13, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i3/graphics/contents/
63hector.nb;   http://www.
mathematica-journal.com/issue/
v6i3/graphics/contents/63hector.
pdf.

Hohmuth:1999:MTR

[Hoh99] Lars Hohmuth. Mathematica 4.2:

a technical review. Mathematica Journal, 9(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Horwitz:2005:MHW

[Hor05] Alan Horwitz. Making holes and windows in surfaces. Mathematica Journal, 10(4):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Hill:1991:EFL

[HR92] L. Thomas Hill and Clifford A. Reiter. Education forum: Laboratories as an enhancement to calculus. Mathematica Journal, 2 (1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/columns/forum/
                                        index.html.

Hrabovsky:2012:DMF

[Hra12] George E. Hrabovsky. Diffusion modeling: Fourier-space representation of nonlinear dynamics of continuous media in cylindrical geometry exploring singularities of the second kind programming in multiple paradigms. Mathematica Journal, 14(??):??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/2012/03/diffusion-modeling/

.

Hu:1999:EHQ

[Hu99] Yifan Hu. Efficient, high-quality force-directed graph drawing. Mathematica Journal, 10(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Hutchinson:1995:MPM

[Hut95] George Hutchinson. Manipulating polynomials with multiple variables. Mathematica Journal, 5(3):??, Summer 1995. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i3/article/hutchinson/
index.html.

Jacobson:1992:ETF

[Jac92a] David Jacobson. Engineer’s toolbox: Floating point in Mathematica. Mathematica Journal, 2(3): ??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i3/tutorials/toolbox/
index.html.

Jacobson:1992:ETH

[Jac92b] David Jacobson. Engineer’s toolbox: Hashing in Mathematica. Mathematica Journal, 2(4): ??, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i4/tutorials/toolbox/
index.html.

Jacobson:1991:ETI

[Jac92c] David Jacobson. The engineer’s toolbox: Improving the performance of your programs. Mathematica Journal, 1(4):??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i4/columns/toolbox/
index.html.

Jacobson:1991:ETB

[Jac92d] David Jacobson. The engineer’s toolbox: Of blocks and pipes. Mathematica Journal, 2 (1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/columns/toolbox/
                                        index.html.

Jakschewitz:1994:H

[Jak94] Roland Jakschewitz. Holaphics. Mathematica Journal, 4 (3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Jakschewitz:1996:RE

[Jak96] Roland Jakschewitz. Rendering experiments. Mathematica Journal, 6(4):6–9, Fall 1996. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v6i4/graphics/
                                        contents/rendering.pdf.

Jankowski:1998:LMI

[Jan98] Mariusz Jankowski. Linear and morphological image processing. Mathematica Journal, 8(1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Jirstrand:1998:CGS

[JG98] Mats Jirstrand and Johan Gunnarson. Code generation for simulation and control applications. Mathematica Journal, 8 (2):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Johannesen:2006:BNP

[Joh06] Ivar G. Johannesen. The Buffon Needle Problem revisited in a pedagogical perspective. Mathematica Journal, 11(2):284–299, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v11i2/contents/Johannesen/
Johannesen.pdf;   http://www.
mathematica-journal.com/issue/
v11i2/Johannesen.html.

Jones:1992:PMS

[Jon92a] Kevin Jones. Profiles in Mathematica: Scott Kim. Mathematica Journal, 2(3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i3/profile/index.
html.

Jones:1992:TM

[Jon92b] Kevin Jones. Teaching with Mathematica. Mathematica Journal, 2(4):??, Fall 1992. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v2i4/profile/
index.html.

Jonassen:2006:MIN

[Jon06] T. M. Jonassen. A Mathematica implementation of nonlinear dynamical systems theory via the spider algorithm and finding critical zeros of high-degree polynomials. Mathematica Journal, 11 (3):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

staff:1991:MOL

[Jou92a] Journal staff. Mathematica oneliners. Mathematica Journal, 1 (4):??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i4/tutorials/oneliners/
                                        index.html.

staff:1991:TT

[Jou92b] Journal staff. Tricks of the trade. Mathematica Journal, 1 (4):??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i4/tutorials/tricks/
                                        index.html.

Jensen:1998:MRF

[JR98] Erik Jensen and Ronald Rusay. Musical representations of the Fibonacci string and proteins. Mathematica Journal, 8(2):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Kabai:1998:CI

[Kab98] Sándor Kabai. Cover image. Mathematica Journal, 8(3):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Kampas:2005:TUR

[Kam05] F. J. Kampas. Tricks of using Reduce to solve Kuhn– Tucker equations. Mathematica Journal, 9(4):686–689,

2005. URL www.mathematica-
journal.com/issue/v9i4/contents/
Tricks9-4/Tricks9-4_2.html.

See correction [Par14].

Kabai:2012:CPP

[KBS12] Sándor Kabai, Szaniszló Bérczi, and Lajos Szilassi. Clusters produced by placing rhombic triacontahedra at the vertices of polyhedra. Mathematica Journal, 14 (??):??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2012/09/clusters-produced-
by-placing-rhombic-triacontahedra-
at-the-vertices-of-polyhedra/

.

Keiper:1991:NC

[Kei92] Jerry Keiper. Numerical computations. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Keiper:1995:IAM

[Kei95] Jerry B. Keiper. Interval arithmetic in Mathematica. Mathematica Journal, 5(2):??, Spring 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (elec-

tronic).  URL http://www.
mathematica-journal.com/issue/
v5i2/article/keiper/index.
html.

Kelly:2006:EFO

[Kel06] Michael Kelly. Evaluation of financial options using radial basis functions in Mathematica. Mathematica Journal, 11(3):??, ????

2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Kendall:1999:SIT

[Ken99] Wilfrid S. Kendall. Stochastic integrals and their expectations. Mathematica Journal, 9 (4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Keranen:1998:IPT

[Ker98] Veikko Keränen. Introduction to Proceedings of the Third International Mathematica Symposium (Hagenberg, Austria, August 23– 25, 1999). Mathematica Journal, 8(2):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Keranen:2006:CW

[Ker06] Veikko Keränen. Combinatorics on words. Mathematica Journal, 11(3):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Kerckhove:2012:PDP

[Ker12] Michael Kerckhove. From population dynamics to partial differential equations. Mathematica Journal, 14(??):??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/2012/04/from-population-
                                        dynamics-to-partial-differential-
                                        equations/.

Khera:1990:IAE

[KG90] Vivek Khera and Henry Greenside. An interface for accessing external numerical libraries. Mathematica Journal, 1(2):84– 88, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/article/khera/
index.html.

Kivela:2006:VRS

[Kiv06] Simo Kivelä. On the visualization of Riemann surfaces. Mathematica Journal, 11(3):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Khemka:2006:ETE

[KJ06] Namrata Khemka and Christian Jacob. Exploratory toolkit for evolutionary and swarm-based optimization. Mathematica Journal, 11(3):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Knopfmacher:1999:OUF

[KM99] Arnold Knopfmacher and Michael Mays. Ordered and unordered factorizations of integers. Mathematica Journal, 10(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Kijewski:1990:VHP

[KMRC90] Marie Foley Kijewski, Stefan P.

Mueller, Frank J. Rybicki, and Paulo A. Carvalho. Very high-precision inversion of illconditioned matrices: a medical imaging application. Mathematica Journal, 1(1):69–74, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i1/features/foley/
index.html.

Knapp:1996:PDE

[Kna96] Robert Knapp. Partial differential equations. Mathematica Journal, 6(4):??, Fall 1996. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v6i4/version3.
0/knapp/index.html.

Knapp:1997:PAM

[Kna97] Rob Knapp. Packed arrays in Mathematica 4. Mathematica Journal, 7(4):??, ???? 1997. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Knuth:1992:CP

[Knu92]  Donald E. Knuth.  Convo-

lution polynomials. Mathematica Journal, 2(4):67–78, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i4/article/knuth/
index.html.

Koepf:1994:PFP

[Koe94]  Wolfram Koepf.  A package

on formal power series. Mathematica Journal, 4(2):62–69, Spring 1994. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i2/article/koepf/
index.html.

Korsan:1993:DUAc

[Kor93a] Robert J. Korsan. Decisions, uncertainty, and all that: An example of Bayesian inference. Mathematica Journal, 3(3):??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/columns/korsan/
index.html.

Korsan:1993:DUAd

[Kor93b] Robert J. Korsan. Decisions, uncertainty, and all that: Flipping thumbtacks. Mathematica Journal, 3(4):??, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i4/columns/korsan/
index.html.

Korsan:1993:DUAa

[Kor93c] Robert J. Korsan. Decisions, uncertainty, and all that: Fractals and time series analysis. Mathematica Journal, 3(1):??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i1/columns/korsan/
index.html.

Korsan:1993:DUAb

[Kor93d] Robert J. Korsan. Decisions, uncertainty, and all that: Mixed-integer linear programming. Mathematica Journal, 3(2): ??, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v3i2/columns/korsan/
                                        index.html.

Korsan:1994:DUAa

[Kor94a] Robert J. Korsan. Decisions, uncertainty, and all that: Nothing ventured, nothing gained: Modeling venture capital decisions. Mathematica Journal, 4(1):74– 80, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i1/columns/korsan/
                                        74-80_Bobk.mj.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v4i1/columns/korsan/
                                        index.html.

Korsan:1994:DUAb

[Kor94b] Robert J. Korsan. Decisions, uncertainty, and all that: Simulate — if you must! Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i3/columns/korsan/
                                        index.html.

Kozlowski:1999:APM

[Koz99] Andrzej Kozlowski. Algebraic programming in Mathematica.

Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Kampas:1999:CAD

[KP99] Frank J. Kampas and János D.

Pintér. Configuration analysis and design by using optimization tools in Mathematica. Mathematica Journal, 10(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Kuchel:1999:SIR

[Kuc99]  Philip W. Kuchel.  Spatial

inversion: Reflective anamorphograms. Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Kuzniarek:1994:GDM

[Kuz94] André Kuzniarek. The graphics designer: Making the perfect picture. Mathematica Journal, 4 (4):??, Fall 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i4/columns/kuzniarek/
index.html.

Kuzniarek:1995:GDG

[Kuz95a] Andre Kuzniarek. The graphic designer: Graphics editing outside of Mathematica. Mathematica Journal, 5(4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v5i4/columns/
andre/index.html.

Kuzniarek:1995:GDP

[Kuz95b] André Kuzniarek. The graphics designer: Putting pictures elsewhere. Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i1/columns/kuzniarek/
                                        index.html.

Levy:1991:TT

[LA92] Silvio Levy and Paul C. Abbott. Tricks of the trade. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i3/tutorials/tricks/
                                        index.html.

Levy:1991:PC

[Lev92a] Silvio Levy. 1992 programming competition. Mathematica Journal, 1(3):??, Winter 1992. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Levy:1992:LLM

[Lev92b] Silvio Levy. L^{A}TEX labels in Mathematica graphics. Mathematica Journal, 2(4):65–66, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i4/article/levy/
                                        index.html.

Levy:1991:OM

[Lev92c] Silvio Levy. Overview of Mathematica 2.0. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Levy:1991:PA

[Lev92d] Silvio Levy. Program announcements. Mathematica Journal, 1 (4):??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i4/columns/announcements/
index.html.

Levy:1996:RG

[Lev96] Silvio Levy. The regal ghost.

Mathematica Journal, 6(1):??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i1/graphics/levy/
index.html.

Lichtblau:1996:GBM

[Lic96] Daniel Lichtblau. Gröbner bases in Mathematica 3.0. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v6i4/article/
lichtblau/index.html.

Liu:1996:C

[Liu96] Bo Liu. The compiler. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v6i4/version3.
                                        0/liu/index.html.

Levy:1990:AE

[LO90] Silvio Levy and Toby Orloff. Automatic Escher. Mathematica Journal, 1(1):??, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Lutovac:1999:ERF

[LT99] Miroslav D. Lutovac and Dejan V. Tosic. Elliptic rational functions. Mathematica Journal, 9 (3):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Lyasoff:1999:PIM

[Lya99]  Andrew Lyasoff.  Path inte-

gral methods for parabolic partial differential equations with examples from computational finance. Mathematica Journal, 9 (2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Lyasoff:2005:DII

[Lya05] Andrew Lyasoff. Dynamic integration of interpolating functions and some concrete optimal stopping problems. Mathematica Journal, 10(4):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Maeder:1990:MPC

[Mae90] Roman E. Maeder. The Mathematica programmer: Computations with infinite structures.

Mathematica Journal, 1(2):30– 33, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/tutorials/maeder/
index.html.

Maeder:1991:SSV

[Mae92a] Roman Maeder. Scoping of symbols and values. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Maeder:1992:MPA

[Mae92b] Roman E. Maeder. The Mathematica programmer: Abstract data types. Mathematica Journal, 2(3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i3/tutorials/maeder/
index.html.

Maeder:1991:MPFa

[Mae92c] Roman E. Maeder. The Mathematica programmer: Fibonacci on the fast track. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i3/tutorials/maeder/
index.html.

Maeder:1991:MPFb

[Mae92d] Roman E. Maeder. The Mathematica programmer: Fractal curves. Mathematica Journal, 1 (4):??, Spring 1992. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i4/columns/maeder/
                                        index.html.

Maeder:1991:MPM

[Mae92e] Roman E. Maeder. The Mathematica programmer: Minimal surfaces. Mathematica Journal, 2(2):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i2/tutorials/maeder/
                                        index.html.

Maeder:1992:MPP

[Mae92f] Roman E. Maeder. The Mathematica programmer: Polymorphism and message passing. Mathematica Journal, 2(4): ??, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i4/tutorials/maeder/
                                        index.html.

Maeder:1991:MPS

[Mae92g] Roman E. Maeder. The Mathematica programmer: System programming. Mathematica Journal, 2(1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/columns/maeder/
                                        index.html.

Maeder:1993:MPO

[Mae93a] Roman E. Maeder. The Mathematica programmer: Objectoriented programming. Mathematica Journal, 3(1):??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i1/columns/maeder/
index.html.

Maeder:1993:MPR

[Mae93b] Roman E. Maeder. The Mathematica programmer: Relational databases. Mathematica Journal, 3(2):??, Spring 1993. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v3i2/columns/
maeder/index.html.

Maeder:1993:MPT

[Mae93c] Roman E. Maeder. The Mathematica programmer: Turing machines and code optimization. Mathematica Journal, 3(3):??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/columns/maeder/
index.html.

Maeder:1993:MPU

[Mae93d] Roman E. Maeder. The Mathematica programmer: Uniform polyhedra. Mathematica Journal, 3(4):??, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i4/columns/maeder/
index.html.

Maeder:1994:MPA

[Mae94a] Roman E. Maeder. The Mathematica programmer: Animated algorithms. Mathematica Journal, 4(4):37–43, Fall 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i4/columns/maeder/
index.html;   http://www.mathematica-
journal.com/issue/v4i4/columns/
maeder/v4i4p37-43.pdf.

Maeder:1994:MPLa

[Mae94b] Roman E. Maeder. The Mathematica programmer: Logic programming I: The interpreter. Mathematica Journal, 4(1):53– 63, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i1/columns/maeder/
53-63_Roman41.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i1/columns/maeder/
index.html.

Maeder:1994:MPLb

[Mae94c] Roman E. Maeder. The Mathematica programmer: Logic programming II: Applications. Mathematica Journal, 4(2):38– 43, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i2/columns/maeder/
38-43_maeder42.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i2/columns/maeder/
index.html.

Maeder:1994:MPR

[Mae94d] Roman E. Maeder. The Mathematica programmer: Ray tracing and graphics extensions. Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i3/columns/maeder/
index.html.

Maeder:1995:BRB

[Mae95a] Roman E. Maeder. Book review: BUGS in writing: Debugging your prose. Mathematica Journal, 5(4):20–21, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/reviews/maeder/
20-21bugs.mj.pdf;   http://
www.mathematica-journal.com/
issue/v5i4/reviews/maeder/
index.html.

Maeder:1995:MPC

[Mae95b] Roman E. Maeder. The Mathematica programmer: Combinators. Mathematica Journal, 5 (4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/columns/maeder/
index.html.

Maeder:1995:MPF

[Mae95c] Roman E. Maeder. The Mathematica programmer: Function iteration and chaos. Mathematica Journal, 5(2):??, Spring 1995.

CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v5i2/columns/
                                        maeder/index.html.

Maeder:1995:MPH

[Mae95d] Roman E. Maeder. The Mathematica programmer: Higherorder functions. Mathematica Journal, 5(3):??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i3/columns/maeder/
                                        index.html.

Maeder:1995:MPS

[Mae95e] Roman E. Maeder. The Mathematica programmer: Singleimage stereograms. Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i1/columns/maeder/
                                        index.html.

Maeder:1996:MPF

[Mae96a] Roman E. Maeder. The Mathematica programmer: Fractional Brownian motion. Mathematica Journal, 6(1):??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i1/columns/maeder/
                                        index.html.

Maeder:1996:MPLa

[Mae96b] Roman E. Maeder. The Mathematica programmer: Long integers: Basic arithmetic and data types. Mathematica Journal, 6 (2):32–40, Spring 1996. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v6i2/columns/
maeder/index.html;   http:
//www.mathematica-journal.
com/issue/v6i2/columns/maeder/
maeder62.pdf.

Maeder:1996:MPLb

[Mae96c] Roman E. Maeder. The Mathematica programmer: Long integers: Efficient algorithms. Mathematica Journal, 6(3):37– 43, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i3/columns/maeder/
contents/63maeder.pdf;   http:
//www.mathematica-journal.
com/issue/v6i3/columns/maeder/
index.html.

Maeder:1996:PF

[Mae96d] Roman E. Maeder. Programming features. Mathematica Journal, 6 (4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i4/version3.0/
maeder/index.html.

Maeder:1997:IPG

[Mae97a] Roman Maeder. Interval plotting and global optimization. Mathematica Journal, 7(3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Maeder:1997:MPP

[Mae97b] Roman Maeder. The Mathematica programmer: Part two: Interval plotting and global optimization. Mathematica Journal, 7(4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Maeder:1997:MPE

[Mae97c] Roman E. Maeder. The Mathematica programmer: Exception handling. Mathematica Journal, 7(1):??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v7i1/columns/maeder/
                                        index.html.

Maeder:1998:FCE

[Mae98] Roman Maeder. Fast computation of equivalence classes. Mathematica Journal, 8(1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Malacarne:2014:CCA

[Mal14] Rodrigo Loureiro Malacarne. Canonical correlation analysis. Mathematica Journal, 16(??):??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/2014/06/canonical-correlation-
                                        analysis/.

Martin:1998:ECK

[Mar98] Scot T. Martin. Employing complex kinetic diagrams. Mathematica Journal, 8(1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Markowsky:2011:TBR

[Mar11] Greg Markowsky. Two basic results concerning random walks on graphs. Mathematica Journal, 13 (??):??, ???? 2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2011/09/two-basic-results-
concerning-random-walks-on-
graphs/.

Mathews:1995:BRI

[Mat95] Kirk A. Mathews. Book review: Introduction to Programming with Mathematica: Mathematica programming — start here. Mathematica Journal, 5(4): 17–19, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/reviews/mathews/
17-19mathews.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v5i4/reviews/mathews/

index.html. Introduction to Programming with Mathematica, by Richard J. Gaylord, Samuel N. Kamin, and Paul R. Wellin. TELOS/ Springer-Verlag, 1993. 302 pages, hardcover, with disk. $39.95. ISBN 0-387-94048-0. Second edition to be published in September 1995.

Matiyasevich:1998:HCG

[Mat98] Yuri Matiyasevich. How to color a graph with (computer) algebra. Mathematica Journal, 8 (4):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

McClure:2006:GSA

[McC06] Mark McClure. Generating selfaffine tiles and their boundaries. Mathematica Journal, 11(1):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

McPherson:1992:MBF

[McP92a] A. McPherson. Modal behavior of the four-layer planar waveguide. Mathematica Journal, 2(1):75–77, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

McPherson:1991:MBF

[McP92b] Angela L. McPherson. Modal behavior of the four-layer planar waveguide. Mathematica Journal, 2(1):75–77, ???? 1992. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v2i1/article/
                                        mcpherson/index.html.

Malsch:1999:ACS

[MD99] Elisabeth Anna Malsch and Gautam Dasgupta. Algebraic construction of smooth interpolants on polygonal domains. Mathematica Journal, 9(3):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Mezzino:1992:ITE

[Mez92] Michael J. Mezzino, Jr. Including TEX expressions in a notebook. Mathematica Journal, 2(4): 61–64, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i4/article/mezzino/
index.html.

Mikhailov:2012:WF

[MF12] Mikhail D. Mikhailov and Atila P. Silva Freire. The Walker function. Mathematica Journal, 14 (??):??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2012/06/the-walker-function/

.

Matis:2010:AMC

[MG10] Timothy I. Matis and Ivan G.

Guardiola. Achieving moment closure through cumulant neglect. Mathematica Journal, 12(1):2:1– 2:18, ???? 2010. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/data/uploads/2010/10/Matis.
pdf.

Murrell:2011:FDK

[MHT11] Hugh Murrell, Kazuo Hashimoto, and Daichi Takatori. Fisher discrimination with kernels. Mathematica Journal, 13(??):??, ???? 2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.

                                        com/2011/07/fisher-discrimination-
                                        with-kernels/.

Miki:2013:ULM

[Mik13] Yasushi Miki. Using the logistic map to generate scratching sounds. Mathematica Journal, 15(??):??, ???? 2013. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/2013/05/using-
                                        the-logistic-map-to-generate-
                                        scratching-sounds/.

Miller:1990:CAF

[Mil90]  Ross M. Miller.  Computer-

aided financial analysis: An implementation of the Black– Scholes model. Mathematica Journal, 1(1):75–79, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i1/features/miller/
                                        index.html.

Miller:1997:TSC

[Mil97] Bruce Miller. Technical support: Customizing buttons and dialogs. Mathematica Journal, 7(3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Marin:1998:CDC

[MIS98] Mirceqa Marin, Tetsuo Ida, and Wolfgang Schreiner. CFLP: a distributed constraint system. Mathematica Journal, 8(2):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Mitic:1996:CVW

[Mit96] Peter Mitic. Critical values for the Wilcoxon signed rank statistic. Mathematica Journal, 6(3):73– 77, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i3/article/mitic/
contents/63mitic.pdf;   http:
//www.mathematica-journal.
com/issue/v6i3/article/mitic/
index.html.

Moore:1993:RIS

[MJ93] Ross Moore and Scott Johnson.

Reconstructing inclusion surfaces within metamorphic garnet crystals. Mathematica Journal, 3(2): ??, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i2/article/moore/
index.html.

Mitchell:1993:BLS

[MM93] C. E. J. Mitchell and A. B.

McLean. Bravais lattices, surface nets, and Buckminsterfullerenes. Mathematica Journal, 3(3):65– 68, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/article/mitchell/
index.html.

Miyadera:2013:CGH

[MNOH13] Ryohei Miyadera, Shunsuke Nakamura, Yu Okada, and Ryo Hanafusa. Chocolate games: How high school students discovered new formulas using Mathematica. Mathematica Journal, 15(??): ??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2013/12/chocolate-games/

.

Moore:1995:SS

[Moo95] Ross Moore. Slicing solids. Mathematica Journal, 5(4):8– 11, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/graphics/moore/
08-11moore.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v5i4/graphics/moore/
index.html.

Murrell:1992:ARR

[Mur92] Hugh Murrell. Animation of rotating rigid bodies. Mathematica Journal, 2(1):62–65, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i1/article/murrell/
index.html.

Murrell:1993:MGS

[Mur93] Hugh Murrell. A mathematical golf swing. Mathematica Journal, 3(4):62–65, Fall 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i4/article/murrell/
index.html.

Murrell:1994:PPP

[Mur94] Hugh Murrell. Planar phase plots and bifurcation animations. Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i3/article/murrell/
index.html.

Murrell:1996:CAT

[Mur96] Hugh Murrell. Computeraided tomography. Mathematica Journal, 6(2):60–65, Spring 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i2/article/murrell/
index.html;   http://www.mathematica-
journal.com/issue/v6i2/article/
murrell/murrell.pdf.

Muresan:2012:SLM

[Mur12] Marian Muresan. Soft landing on the Moon with Mathematica. Mathematica Journal, 14(??): ??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2012/08/soft-landing-on-
the-moon-with-mathematica/.

Muhandiramge:2005:ESM

[MW05] Ranga D. Muhandiramge and Jingbo Wang. Electronic structure of multi-electron quantum dots. Mathematica Journal, 10 (2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Nachbar:1995:GP

[Nac95]  Robert B. Nachbar.  Ge-

netic programming. Mathematica Journal, 5(3):44–55, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v5i3/tutorials/nachbar/
                                        index.html.

Nagasaka:2005:SNA

[Nag05] Kosaku Nagasaka. Symbolicnumeric algebra for polynomials. Mathematica Journal, 10 (3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Nakos:1993:EPM

[Nak93]  George Nakos.  Expansions

of powers of multivariate formal power series. Mathematica Journal, 3(1):??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v3i1/article/nakos/
                                        index.html.

Nilsson:1999:FIS

[NBT99] Roland Nilsson, Johan Björkegren, and Jesper Tegnér. A flexible implementation for support vector machines. Mathematica Journal, 10(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Nerlove:1999:NAM

[Ner99] Marc Nerlove. On the numerical accuracy of Mathematica 5.0 for doing linear and nonlinear regression. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Neves:1999:ASL

[Nev99] Armando G. M. Neves. Approximating solutions of linear ordinary differential equations with periodic coefficients by exact Picard iterates. Mathematica Journal, 10(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Nakos:1994:CGB

[NG94] George Nakos and Nikolaos Glinos. Computing Gröbner bases over the integers. Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i3/article/nakos/
index.html.

Noguchi:1991:TGE

[Nog92a] A. Noguchi. The two-sector general equilibrium model. Mathematica Journal, 1(3):96–103, Winter 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Noguchi:1991:TSG

[Nog92b] Asahi Noguchi. Two-sector general equilibrium model: Numerical and graphical representations of an economy. Mathematica Journal, 1(3):96–103, Winter 1992. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i3/article/noguchi/
index.html.

Norton:1995:SE

[Nor95] Andrew H. Norton. Spinors and entanglement. Mathematica Journal, 5(2):??, Spring 1995. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v5i2/graphics/
norton/index.html.

Notebook:1996:IMP

[Not96] Christian Hidber Notebook. Integrating Mathematica into Prolog. Mathematica Journal, 6(3): ??, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i3/article/hidber/
contents/63prolog.nb;   http:
//www.mathematica-journal.
com/issue/v6i3/article/hidber/
index.html.

Novak:1993:W

[Nov93] John M. Novak. WorldPlot. Mathematica Journal, 3(3):??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/graphics/index.
html.

Novak:1996:BP

[Nov96a] John M. Novak. Buttons and palettes. Mathematica Journal, 6 (4):??, Fall 1996. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i4/version3.0/
novak/index.html.

Novak:1996:PN

[Nov96b] John M. Novak. Programming notebooks. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i4/version3.0/
novak2/index.html.

Novak:1999:WEW

[Nov99] John M. Novak. WaveX: Extracting wavelets from seismic data. Mathematica Journal, 9 (3):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Nethery:1995:MBF

[NS95] John F. Nethery and Mark W.

Spong. A MathLink-based front end for Robotica package. Mathematica Journal, 5(2): ??, Spring 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i2/article/nethery/
index.html.

Ogawa:1991:H

[Oga92a] Arthur Ogawa. Helicatenoid.

Mathematica Journal, 2(2):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.

                                        com/issue/v2i2/graphics/index.
                                        html.

Ogawa:1991:TR

[Oga92b] Arthur Ogawa. The trinoid revisited. Mathematica Journal, 2 (1):??, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/graphics/ogawa/
                                        index.html.

Palancz:2014:FDD

[Pal14]  Béla Paláncz.  Fitting data

with different error models. Mathematica Journal, 16(??):??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/2014/04/fitting-data-with-
                                        different-error-models/.

Papegay:2005:EBG

[Pap05] Yves Papegay. Exploring board game strategies. Mathematica Journal, 10(2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Parreiras:2014:URC

[Par14] Sérgio O. Parreiras. Using Reduce to compute Nash equilibria: Classroom tools for game theory. Mathematica Journal, 16(??):??, ???? 2014. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/2014/03/using-
                                        reduce-to-compute-nash-equilibria/

. See [Kam05].

Murphy:1999:MG

[PC99] Paul R. McCreary, Teri Jo Murphy and Christan Carter. The modular group. Mathematica Journal, 9(3):??, ???? 1999. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Pegg:2005:BSC

[Peg05a] Ed Pegg, Jr. Beyond sudoku: Classic puzzles in Wolfram demonstrations. Mathematica Journal, 10(4):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Pegg:2005:BSN

[Peg05b] Ed Pegg, Jr. Beyond sudoku:

Number link. Mathematica Journal, 10(3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Pegg:2005:BSR

[Peg05c] Ed Pegg, Jr. Beyond sudoku:

Ripple effect. Mathematica Journal, 10(2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Pegg:2006:BSC

[Peg06a] Ed Pegg, Jr. Beyond sudoku:

Crossing number graphs. Mathematica Journal, 11(2):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Anonymous:2006:C

[Peg06b] Ed Pegg, Jr. Beyond sudoku: The Icosian game, revisited. Mathematica Journal, 11(3):??, ????

2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Petersen:1991:DE

[Pet92]  Troels Petersen.  Differential

equations. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Petersen:1996:OM

[Pet96] Troels Petersen. Overview of Mathematica 3.0. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v6i4/version3.
0/petersen/index.html.

Papusha:2006:SEB

[PFS06] Alexander N. Papusha, Ilya V. Fedorov, and Victor V. Shtrasser. Symbolic evaluation of boundary problems for offshore design technology. Mathematica Journal, 11 (1):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Phillips:1990:UMI

[Phi90] Mark B. Phillips. Using Mathematica as an interface to an interactive graphics program. Mathematica Journal, 1(2):45– 50, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/article/phillips/
index.html.

Phillips:1997:EPA

[Phi97a] Flip Phillips. Editor’s pick: Artlandia: Cool tools for computer art. Mathematica Journal, 7 (3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Phillips:1997:SM

[Phi97b] Flip Phillips. SimCity 3000 and Mathematica. Mathematica Journal, 7(4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Phillips:1998:EPQ

[Phi98a] Flip Phillips. Editor’s pick:

Quantum computation. Mathematica Journal, 8(1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Phillips:1998:LUR

[Phi98b] Flip Phillips. Limits, uncertainty, and randomness. Mathematica Journal, 8(3):??, ???? 1998. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Provost:1999:MBD

[Pro99] Serge B. Provost. Moment-based density approximants. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Piziak:1998:RQR

[PT98] Robert Piziak and Danny W.

Turner. Real quaternions and rotations. Mathematica Journal, 8 (3):??, ???? 1998. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v8i3/features/turner/
contents/html/index.html;
http://www.mathematica-journal.
com/issue/v8i3/features/turner/
contents/notebooks/Quaternions.
nb.

Porta:1991:EFT

[PUS92] Horacio Porta, Jerry Uhl, and Francis Sullivan. Education forum: Three excerpts from euler. Mathematica Journal, 1(4): ??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i4/columns/forum/
index.html.

Ramsden:1997:MCC

[Ram97] Phil Ramsden. Mathematica in the classroom: Classroom chemistry simulations. Mathematica Journal, 7(3):??, ???? 1997. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Reiter:1992:FFN

[Rei92]  Clifford Reiter.  Fast Fi-

bonacci numbers. Mathematica Journal, 2(3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i3/article/reiter/
index.html.

Reinking:1998:OSR

[Rei98] J. Todd Reinking. Orthogonal series representation of probability density and distribution functions. Mathematica Journal, 8 (3):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Richardson:1999:IE

[Ric99] Stan Richardson. Integral equations. Mathematica Journal, 9 (2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Riddle:1990:NCA

[Rid90]  Alfy Riddle.  A nodal cir-

cuit analysis program. Mathematica Journal, 1(1):62–68, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i1/features/riddle/
index.html.

Rubenkonig:2005:IL

[RK05] Oliver Rübenkönig and Jan G.

Korvink. Interactive learning. Mathematica Journal, 10(3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Rubenkonig:2005:IED

[RLK05] Oliver Rübenkönig, Zhenyu Liu, and Jan G. Korvink. Integrated engineering development environment. Mathematica Journal, 10 (3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Rangel-Mondrag:1999:PRF

[RM99] Jaime Rangel-Mondragón. Polyominoes and related families. Mathematica Journal, 9(3):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Rangel-Mondragon:2006:BJP

[RM06] Jaime Rangel-Mondragón. Buckets and jumping pets. Mathematica Journal, 11(1):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Rangel-Mondragon:2013:STC

[RM13] Jaime Rangel-Mondragon. Selected themes in computational non-Euclidean geometry: Part 1: Properties of inversive geometry. Mathematica Journal, 15 (??):??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2013/07/selected-themes-
in-computational-non-euclidean-
geometry-part-1/.

Rangel-Mondragon:2014:AFN

[RM14] Jaime Rangel-Mondragón. The arbelos: Incircle, radical circle, radical axis, twins, generalizations, and proofs without words. Mathematica Journal, 16 (??):??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2014/05/the-arbelos/.

Riste:2005:AWP

[RMWP05] Neil Riste, Bradley McGrath, Jingbo Wang, and Jie Pan. Acoustic wave propagator — A split region implementation. Mathematica Journal, 10(3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Rimmer:1999:SDM

[RN99] Robert H. Rimmer and John P.

Nolan. Stable distributions in Mathematica. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Roberts:2005:PED

[RN05] Dale Olivier Roberts and Alexander Novikov. Pricing European and discretely monitored exotic options under the Lévy Process Framework. Mathematica Journal, 10(3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Robinson:1995:W

[RR95] Sam L. Robinson and Peter F. Ryczek. Wavelets. Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i1/article/robinson/
index.html.

Ramirez:2014:PGF

[RR14] José L. Ramírez and Gustavo N.

Rubiano. Properties and generalizations of the Fibonacci word fractal: Exploring fractal curves. Mathematica Journal, 16(??):??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2014/02/properties-and-
generalizations-of-the-fibonacci-
word-fractal/.

Rose:1996:MND

[RS96] Colin Rose and Murray D. Smith. The multivariate normal distribution. Mathematica Journal, 6(1): ??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i1/tutorials/rose/
index.html.

Rose:1999:COS

[RS99] Colin Rose and Murray D. Smith. Computational order statistics. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Rudd:1994:VS

[Rud94] Robert Rudd. Virtual slinky. Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i3/graphics/rudd/
index.html.

Rudd:1995:LL

[Rud95a] Robert Rudd. Lava lamp. Mathematica Journal, 5(4):12– 15, Fall 1995. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/graphics/rudd/
12-16rudd.mj.pdf;   http://
www.mathematica-journal.com/
issue/v5i4/graphics/rudd/index.
html.

Rudd:1995:MAP

[Rud95b] Robert Rudd. A Mathematica amusement park. Mathematica Journal, 5(1):??, Winter 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i1/graphics/index.
html.

Rudd:1995:MD

[Rud95c] Robert Rudd. Möbius dominoes. Mathematica Journal, 5(2): ??, Spring 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i2/graphics/rudd/
index.html.

Rudd:1997:W

[Rud97] Robert Rudd. Waterworks. Mathematica Journal, 7(3):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Rudd:1998:LR

[Rud98] Robert Rudd. Luge ride. Mathematica Journal, 8(1):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Rudd:2006:EMS

[Rud06] Robert Rudd. Estimating the mu slip curve via extended Kalman filtering. Mathematica Journal, 11(1):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Smith:1990:MP

[SA90] Cameron Smith and Paul Abbott. Manipulating polygons. Mathematica Journal, 1(1):??, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Saarinen:1996:NIO

[Saa96]  Sirpa Saarinen.  Numerical

integration and optimization. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i4/version3.0/
                                        saarinen/index.html.

Sandri:1996:NCL

[San96] Marco Sandri. Numerical calculation of Lyapunov exponents. Mathematica Journal, 6(3):78– 84, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i3/article/sandri/
                                        contents/63sandri.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v6i3/article/sandri/
                                        index.html.

Sandhu:1999:APD

[San99a] Pavi Sandhu. AuthorTools: a package for document processing. Mathematica Journal, 9 (2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Sandhu:1999:CHB

[San99b] Pavi Sandhu. Customizing the help browser. Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Sandhu:1999:XM

[San99c] Pavi Sandhu. XML and Mathematica. Mathematica Journal, 9(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Sarafian:1999:SSM

[Sar99] Haiduke Sarafian. The second set of magic angles of projectile motion. Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Sarafian:2011:SSN

[Sar11] Haiduke Sarafian. A study of super-nonlinear motion of a simple pendulum. Mathematica Journal, 13(??):??, ???? 2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/2011/08/a-study-
of-super-nonlinear-motion-
of-a-simple-pendulum/.

Simon:1993:PSR

[SB93] Julian L. Simon and Peter Bruce. Probability and statistics with resampling stats and Mathematica. Mathematica Journal, 3(1): ??, Winter 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i1/article/simon/
index.html.

Smith:1994:LSG

[SB94a] Cameron Smith and Nancy Blachman. Lighting and shading in 3D graphics. Mathematica Journal, 4(3):??, Summer 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i3/tutorials/smith/
index.html.

Smith:1994:PPC

[SB94b] Cameron Smith and Nancy Blachman. Perspectives projection and coordinates in 3D graphics. Mathematica Journal, 4(2): ??, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i2/tutorials/smith/
index.html.

Selig:1994:SRB

[Sel94] J. M. Selig. Some rule-based programming examples. Mathematica Journal, 4(1):44–47, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i1/tutorials/selig/
43-46_Selig.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i1/tutorials/selig/
index.html.

Senechal:1994:TDQ

[Sen94]  Marjorie Senechal.  Tilings,

diffraction, and quasicrystals. Mathematica Journal, 4(2):10– 15, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i2/graphics/10-
15_senechal42.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i2/graphics/index.
html.

Sheshadri:2011:MPS

[SF11] K. Sheshadri and Peter Fritzson. MathPDE: A package to solve PDEs by finite differences. Mathematica Journal, 13(??):??, ???? 2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2011/12/mathpde-a-package-
to-solve-pdes-by-finite-differences/

.

Shaw:1995:SCC

[Sha95] William T. Shaw. Symmetric chaos in the complex plane. Mathematica Journal, 5(3):??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i3/article/shaw/
index.html.

Silvestri:2013:RWW

[Sil13] Todd Silvestri. Random walks on the World Wide Web. Mathematica Journal, 15(??):??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/2013/09/random-
                                        walks-on-the-world-wide-web/

.

Silvestri:2014:CSR

[Sil14] Todd Silvestri. Complex system reliability: A graph theory approach. Mathematica Journal, 16 (??):??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/2014/07/complex-system-
                                        reliability/.

Siniksaran:2006:TBN

[Sin06] Enis Siniksaran. Throwing Buffon’s Needle with Mathematica. Mathematica Journal, 11(1):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Skolnick:1992:ARR

[Sko92] Harold Skolnick. Art has reasons that reason does not understand. Mathematica Journal, 2 (3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Skottun:1994:APP

[Sko94] Brent C. Skottun. On amplitude and phase in printed characters. Mathematica Journal, 4(2):

83–86, Spring 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i2/article/skottun/
83-86_skottun42.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i2/article/skottun/
index.html.

Sato:1998:ATD

[SMH98] Hiroki Sato, Kenichi Muro, and Akira Hasegawa. Animated three-dimensional structure of the Earth’s interior for research and education. Mathematica Journal, 8(4):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v8i4/features/sato/
index.html.

Smith:1990:FD

[Smi90a] Cameron Smith. Facing the data.

Mathematica Journal, 1(2):34– 37, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/tutorials/smith/
index.html.

Smith:1990:OPN

[Smi90b] Cameron Smith. Opening packages as notebooks. Mathematica Journal, 1(1):??, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Smith:1991:G

[Smi92a] Cameron Smith. Graphics. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Smith:1991:IT

[Smi92b] Cameron Smith. Illustrating TEX. Mathematica Journal, 1(3): ??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v1i3/tutorials/smith/
                                        index.html.

Smith:1993:NBL

[Smi93]  Cameron Smith.  Notebooks

into books via L^{A}TEX. Mathematica Journal, 3(3):69–73, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v3i3/article/smith/
                                        index.html.

Saunders:2006:SL

[SN06] Andrew Saunders and Amie Nulman. Surface logic. Mathematica Journal, 11(3):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Storti:1992:TIS

[SNG92] Duane Storti, Cornelius Nevrinceanu, and Mark Ganter. A tutorial on implicit solid modeling. Mathematica Journal, 2(3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i3/article/storti/
                                        index.html.

Sofroniou:1993:EBF

[Sof93] Mark Sofroniou. Extending the built-in format rules. Mathematica Journal, 3(3):74–??, Summer 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i3/article/sofroniou/
index.html.

Sofroniou:1996:NM

[Sof96] Marc Sofroniou. Numerics in Mathematica 3.0. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v6i4/article/
sofroniou/index.html.

Soiffer:1996:MT

[Soi96] Neil Soiffer. Mathematical typesetting. Mathematica Journal, 6 (4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i4/version3.0/
soiffer/index.html.

Scholl:1990:TTM

[SR90] J. F. Scholl and D. Rogovin. Transient two-wave mixing in Kerr media. Mathematica Journal, 1(2):81–83, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/article/scholl/
index.html.

Sullivan:1990:AHE

[ST90] John M. Sullivan and Matthew M. Thomas. Animating the heat equation: a case study in optimization. Mathematica Journal, 1(1):80–84, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v1i1/features/
                                        sullivan/index.html.

Stone:1995:C

[Sto95]  John Stone.  Conchyallome-

try. Mathematica Journal, 5 (3):??, Summer 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Strzebonski:1994:ASS

[Str94] Adan W. Strzebonski. An algorithm for systems of strong polynomial inequalities. Mathematica Journal, 4(4):??, Fall 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v4i4/article/
                                        strzebonski/index.html.

Strzebonski:1996:ANM

[Str96] Adam W. Strzebonski. Algebriac numbers in Mathematica 3.0. Mathematica Journal, 6(4): ??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i4/article/strzebonski/
                                        index.html.

Strzebonski:1997:SAI

[Str97] Adam Strzebonski. Solving algebraic inequalities. Mathematica Journal, 7(4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Sullivan:1991:GRF

[Sul92] John Sullivan. Generating and rendering four-dimensional polytopes. Mathematica Journal, 1(3):??, Winter 1992. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v1i3/article/
sullivan/index.html.

Sutner:1992:FSM

[Sut92a] K. Sutner. Finite state machines and syntactic semigroups. Mathematica Journal, 2(1):78–87, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Sutner:1991:FSM

[Sut92b] Klaus Sutner. Finite state machines and syntactic semigroups. Mathematica Journal, 2(1):78– 87, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v2i1/article/sutner/
index.html.

Sutner:2006:DSC

[Sut06] Klaus Sutner. Divisibility and state complexity. Mathematica Journal, 11(3):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Spada:1990:CAA

[SYS^{+}90] Giorgio Spada, David A. Yuen, Roberto Sabadini, Paul J. Morin, and Paolo Gasperini. A computer-aided, algebraic approach to the post-glacial rebound problem. Mathematica Journal, 1(2):65–69, Fall 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i2/article/spada/
index.html.

Tannenbaum:1997:PWV

[Tan97]  Peter Tannenbaum.  Power

in weighted voting systems. Mathematica Journal, 7(1):??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v7i1/article/tannenbaum/
index.html.

Trzeciakowski:1994:PSP

[TC94] Jerome P. Trzeciakowski and William M. Chilian. Polar symmetry plots. Mathematica Journal, 4(4):??, Fall 1994. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v4i4/article/
trzeciakowski/index.html.

Team:1999:WNM

[The99] The Mathematica 5 Product Team. What’s new in Mathematica 5. Mathematica Journal, 9(1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Thimbleby:1993:GR

[Thi93] Harold Thimbleby. Genaille’s rods. Mathematica Journal, 3(2): ??, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i2/article/thimbleby/
index.html.

Thimbleby:1995:EDF

[Thi95] Harold Thimbleby. An example of distribution-free testing. Mathematica Journal, 5(2):??, Spring 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i2/tutorials/thimbleby/
index.html.

Romeny:2005:DR

[tHR05] Bart M. ter Haar Romeny. Dither removal. Mathematica Journal, 10(2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Takagi:1996:MMG

[TiS96] Hideaki Takagi and Ken ichi Sakamaki. Moments for M/G/1 queues. Mathematica Journal, 6(1):??, Winter 1996. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v6i1/article/
index.html.

Towle:1996:PZ

[Tow96] Russell Towle. Polar zonohedra. Mathematica Journal, 6(2): 8–12, Spring 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i2/graphics/index.
                                        html;   http://www.mathematica-
                                        journal.com/issue/v6i2/graphics/
                                        towle.pdf.

Trott:1996:VES

[Tro96] Michael Trott. The vibrating ellipse-shaped drum. Mathematica Journal, 6(4):??, Fall 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v6i4/article/
                                        trott/index.html.

Trott:1997:TCS

[Tro97a] Michael Trott. Trott’s corner: Solving puzzles with Mathematica. Mathematica Journal, 7(3): ??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:1997:TCV

[Tro97b] Michael Trott. Trott’s corner: Visualization of Riemann surfaces. Mathematica Journal, 7 (4):??, ???? 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:1998:TCCa

[Tro98a] Michael Trott. Trott’s corner: Compositions of elementary transcendental functions. Mathematica Journal, 8(1):??, ???? 1998.

CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:1998:TCCb

[Tro98b] Michael Trott. Trott’s corner:

Compositions of elementary transcendental functions. Mathematica Journal, 8(3):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:1998:TCV

[Tro98c] Michael Trott. Trott’s corner:

Visualization of Riemann surfaces. Mathematica Journal, 8 (4):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:1999:TCMb

[Tro99a] Michael Trott. Trott’s corner:

Mathematical searching of The Wolfram Functions Site. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:1999:TCMa

[Tro99b] Michael Trott. Trott’s corner:

Modular equations of the Rogers– Ramanujan continued fraction. Mathematica Journal, 9(2):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:1999:TCSb

[Tro99c] Michael Trott. Trott’s corner:

Scattering states in a 2D quantum dot embedded in a waveguide. Mathematica Journal, 10 (1):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:1999:TCSa

[Tro99d] Michael Trott. Trott’s corner: The sound of an unusual drum. Mathematica Journal, 9 (3):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:2005:TCF

[Tro05a] Michael Trott. Trott’s corner: Finding Trott constants. Mathematica Journal, 10(2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:2005:TCR

[Tro05b] Michael Trott. Trott’s corner: The return of the Riemann surface. Mathematica Journal, 10 (4):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Trott:2006:TCC

[Tro06]  Michael Trott.  Trott’s cor-

ner: Constructing crossword arrays faster. Mathematica Journal, 11(2):??, ???? 2006. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Vardi:1991:SEP

[Var92]  Ilan Vardi.  A sampler of

elegant programs. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.

com/issue/v1i3/tutorials/vardi/
index.html.

Vardi:1996:CPC

[Var96] Ilan Vardi. Code and pseudo code. Mathematica Journal, 6(2): 66–71, Spring 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i2/article/vardi/
index.html;   http://www.mathematica-
journal.com/issue/v6i2/article/
vardi/vardi.pdf.

Varian:1999:BT

[Var99] Hal Varian. Bootstrap tutorial. Mathematica Journal, 9 (4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

vanGastel:1999:UGD

[vGU99] Leendert van Gastel and Harry Uitermark. Updating a geographic database. Mathematica Journal, 9(2):??, ???? 1999. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Vinod:1999:SDE

[Vin99] H. D. Vinod. Skew densities and ensemble inference for financial economics. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Vrbik:2011:BTP

[Vrb11a] Jan Vrbik. Betting two patterns against each other. Mathematica Journal, 13(??):??, ???? 2011. CODEN ????

ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2011/09/betting-two-patterns-
against-each-other/.

Vrbik:2011:SDM

[Vrb11b] Jan Vrbik. Sampling distribution of ML estimators: Cauchy example. Mathematica Journal, 13(??): ??, ???? 2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2011/12/sampling-distribution-
of-ml-estimators-cauchy-example/

.

Vrbik:2011:SKP

[Vrb11c] Jan Vrbik. Solving Kepler’s problem. Mathematica Journal, 13 (??):??, ???? 2011. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2011/07/solving-keplers-
problem/.

Vrbik:2012:GRF

[Vrb12a] Jan Vrbik. Gambler’s ruin and first passage time. Mathematica Journal, 14(??):??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/2012/03/gamblers-
ruin-and-first-passage-time/

.

Vrbik:2012:RMC

[Vrb12b] Jan Vrbik. Relativistic motion of a charged particle and Pauli algebra. Mathematica Journal, 14 (??):??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2012/05/relativistic-motion-
of-a-charged-particle-and-
pauli-algebra/.

Wagon:1993:HIB

[Wag93] Stan Wagon. A hyperbolic interpretation of the Banach–Tarski paradox. Mathematica Journal, 3(4):58–62, Fall 1993. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v3i4/article/
wagon/index.html.

Wagner:1995:PPD

[Wag95] David B. Wagner. Power programming: Dynamic programming. Mathematica Journal, 5 (4):??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v5i4/columns/wagner/
index.html.

Wagner:1996:PPC

[Wag96a] David B. Wagner. Power programming: Contexts and shadowing. Mathematica Journal, 6(2): 41–51, Spring 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i2/columns/wagner/
index.html;   http://www.mathematica-
journal.com/issue/v6i2/columns/
wagner/wagner62.pdf.

Wagner:1996:PPD

[Wag96b] David B. Wagner. Power programming: Dependency analysis. Mathematica Journal, 6(1): ??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i1/columns/wagner/
                                        index.html.

Wagner:1996:PPM

[Wag96c] David B. Wagner. Power programming: MathLink mode. Mathematica Journal, 6(3):44– 57, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v6i3/columns/wagner/
                                        contents/63wagner.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v6i3/columns/wagner/
                                        index.html.

Wagon:1996:GM

[Wag96d] Stan Wagon. Gaussian moats. Mathematica Journal, 6(1):??, Winter 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Wagner:1997:PPF

[Wag97] David B. Wagner. Power programming: Functional iteration. Mathematica Journal, 7(1):??, Winter 1997. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v7i1/columns/wagner/
                                        index.html.

Watson:1990:ELP

[Wat90]  Mark J. Watson.  An ele-

mentary logic package. Mathematica Journal, 1(1):85–91, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i1/features/watson/
index.html.

Watson:1994:ICU

[Wat94]  Andrew B. Watson.  Im-

age compression using the discrete cosine transform. Mathematica Journal, 4(1):81–88, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v4i1/article/81-
88_Watson.mj.pdf;   http://
www.mathematica-journal.com/
issue/v4i1/article/index.html.

Jacobson:1991:MB

[wDJK92] Steve Christensen with David Jacobson and Robert J. Korsan. From the E-mail bag. Mathematica Journal, 1(4):??, Spring 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v1i4/columns/emailbag/
index.html.

Wei:1992:GD

[Wei92] Sha Xin Wei. Global demographics. Mathematica Journal, 2(4): ??, Fall 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i4/graphics/index.
                                        html.

Wei:1994:BRI

[Wei94] Sha Xin Wei. Book review: InterCall 2.1: The tortoise and the hare. Mathematica Journal, 4(1): ??, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v4i1/reviews/wei/
                                        22-25_xinwei.mj.pdf;   http:
                                        //www.mathematica-journal.
                                        com/issue/v4i1/reviews/wei/
                                        index.html.

Weiss:1999:AGF

[Wei99] Peter Weiß. Applications of generating functions in nonparametric tests. Mathematica Journal, 9(4):??, ???? 1999. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Weisstein:2005:MM

[Wei05] Eric W. Weisstein. Making Math-World. Mathematica Journal, 10 (3):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v10i3/MathWorld.
                                        html.

Wellin:2013:SHH

[Wel13] Paul Wellin. Strings: Chapter 9 programming with Mathematica: An introduction. Mathematica Journal, 15(??):??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/2013/04/strings/

.

Wessels:1993:PSM

[Wes93]  Philip S. Wessels.  Plot-

ting, solving, and mapping. Mathematica Journal, 3(2):??, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v3i2/tutorials/wessels/
index.html.

Winfield:2012:AMO

[Win12] Christopher J. Winfield. Asymptotic methods of ODEs: Fourierspace representation of nonlinear dynamics of continuous media in cylindrical geometry exploring singularities of the second kind. Mathematica Journal, 14 (??):??, ???? 2012. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/2012/03/asymptotic-methods-
of-odes/.

Withoff:1990:PSD

[Wit90] David Withoff. Processing survey data. Mathematica Journal, 1(1):92–94, Summer 1990. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v1i1/features/
withoff/index.html.

Withoff:1991:PC

[Wit92] David Withoff. Program conversion. Mathematica Journal, 1 (3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Withoff:1996:PCF

[Wit96] David Withoff. Procedural control functions. Mathematica Journal, 6(4):??, Fall 1996. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/issue/v6i4/version3.
                                        0/withoff/index.html.

Wickham-Jones:1993:TDC

[WJ93] Tom Wickham-Jones. Threedimensional contour plots. Mathematica Journal, 3(2):??, Spring 1993. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Wolfram:1998:FMM

[Wol98] Stephen Wolfram. Foundations of mathematics and Mathematica. Mathematica Journal, 8(2): ??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Wolfram:2005:FC

[Wol05] Stephen Wolfram. The future of computation. Mathematica Journal, 10(2):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Yelland:2010:ICA

[Yel10] Phillip M. Yelland. An introduction to correspondence analysis. Mathematica Journal, 12(1):4:1– 4:23, ???? 2010. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/data/uploads/2010/10/Yelland.
pdf.

Zeller:1996:TSW

[Zel96]  Tom Zeller.  Tech support:

Warning and error messages. Mathematica Journal, 6(3):58– 64, Summer 1996. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.
com/issue/v6i3/columns/techsupport/
contents/63tech.pdf;   http:
//www.mathematica-journal.
com/issue/v6i3/columns/techsupport/
index.html.

Zeller:1997:TSF

[Zel97] Tom Zeller. Tech support: Frequently asked questions about Version 3.0. Mathematica Journal, 7(1):??, Winter 1997. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

URL http://www.mathematica-
journal.com/issue/v7i1/columns/
techsupport/index.html.

Zizza:1992:APD

[Ziz92a] F. Zizza. Algebraic programming and differential forms. Mathematica Journal, 2(1):88–90, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Zizza:1991:APD

[Ziz92b] Frank Zizza. Algebraic programming and differential forms. Mathematica Journal, 2(1):88– 90, ???? 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

                                        //www.mathematica-journal.
                                        com/issue/v2i1/article/zizza/
                                        index.html.

Zwilling:2013:NBR

[Zwi13] Michael L. Zwilling. Negative binomial regression. Mathematica Journal, 15(??):??, ???? 2013. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

                                        URL http://www.mathematica-
                                        journal.com/2013/06/negative-
                                        binomial-regression/.